### v2

# Matrix systems, algebras, and open maps

Abstract.
Every state on the algebra $\mathrm{M}_{n}$ of complex $n\times n$ matrices restricts to a state on any matrix system. Whereas the restriction to a matrix system is generally not open, we prove that the restriction to every *-subalgebra of $\mathrm{M}_{n}$ is open. This simplifies topology problems in matrix theory and quantum information theory.
Key words and phrases:
quantum state, matrix system, open map, numerical range
2020 Mathematics Subject Classification:
Primary 81P16, 94A15, 46A30; Secondary 47A12, 15A60
In honor of Ilya Matveevich Spitkovsky, for his 70th birthday.
1. Introduction
[…]
$\mathcal{S}(\mathcal{R})$ :=$:=\{\ell\in\mathcal{R}^{\ast}\mid\forall A\in\mathcal{C}(\mathcal{R}):\ell(A)\geq 0,\ell(\mathds{1}_{n})=1\}\,.$
The restriction $\mathcal{S}(\mathrm{M}_{n})\to\mathcal{S}(\mathcal{R})$, $\ell\mapsto\ell|_{\mathcal{R}}$ of states to $\mathcal{R}$ is continuous and affine. Its analytic properties would be perfectly clear if it were not for the openness that fails in Example 1.1. Let $K,L$ be subsets of some Euclidean spaces endowed with their relative topologies [27]. A map $f:K\to L$ is open at $x\in K$ if the image of every neighborhood of $x$ in $K$ is a neighborhood of $f(x)$ in $L$. The map $f$ is open if it is open at every point in $K$.
It is helpful to represent states as matrices. The antilinear isomorphism $\mathcal{R}\to\mathcal{R}^{\ast}$, $A\mapsto\langle A,\,\cdot\,\rangle$ restricts by Lemma 2.2 to the affine isomorphism
$r_{\mathcal{R}}:\mathcal{D}(\mathcal{R})\to\mathcal{S}(\mathcal{R})\,,\quad\rho\mapsto\langle\rho,\,\cdot\,\rangle\,,$
where $\langle A,B\rangle:=\operatorname{tr}(A^{\ast}B)$ is the Frobenius inner product of $A,B\in\mathrm{M}_{n}$,
$\mathcal{H}(\mathcal{R})$ :=$:=\{A\in\mathcal{R}\mid A^{\ast}=A\}\,,$
$\mathcal{C}(\mathcal{R})^{\vee}$ :=$:=\{A\in\mathcal{H}(\mathcal{R})\mid\forall B\in\mathcal{C}(\mathcal{R}):\langle A,B\rangle\geq 0\}\,,$
$\text{and}\quad\mathcal{D}(\mathcal{R})$ :=$:=\{\rho\in\mathcal{C}(\mathcal{R})^{\vee}\mid\operatorname{tr}(\rho)=1\}\,.$
Generalizing a term of von Neumann algebras [10], we refer to the elements of $\mathcal{D}(\mathcal{R})$ as density matrices. The inclusion $\mathcal{C}(\mathcal{R})\subset\mathcal{C}(\mathcal{R})^{\vee}$ can be strict. For example, the density matrix $\operatorname{diag}(-1,5,2)/6$ of the matrix system spanned by $\mathds{1}_{3}$ and $\operatorname{diag}(1,-1,0)$ is indefinite. It is well known, see Rem. 2.8, that the identity $\mathcal{C}(\mathcal{A})=\mathcal{C}(\mathcal{A})^{\vee}$ holds for every *-subalgebra $\mathcal{A}$ of $\mathrm{M}_{n}$.
Example 1.1.
We write block diagonal matrices as direct sums, for instance
$\begin{pmatrix}A&0\\ 0&c\end{pmatrix}=A\oplus c\in\mathrm{M}_{3}\quad\text{for every}\quad A\in\mathrm{M}_{2}\,,\quad c\in\mathbb{C}\cong\mathrm{M}_{1}\,.$
Denoting the imaginary unit by $\mathrm{i}\in\mathbb{C}$ and the Pauli matrices by
$X:=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\,,\quad Y:=\begin{pmatrix}0&-\mathrm{i}\\ \mathrm{i}&0\end{pmatrix}\,,\quad Z:=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\,,$
we define the matrix system $\mathcal{R}:=\operatorname{span}_{\mathbb{C}}(\mathds{1}_{3},X\oplus 1,Z\oplus 0)$. The orthogonal projection of $\mathbb{C}^{2}$ onto the line spanned by |$\ket{+}:=(1,1){}^{\mathrm{T}}/\sqrt{2}$ is written $\ket{+}\!\!\bra{+}$ | in Dirac’s notation [7, 11]. The open set $\mathcal{O}:=\{\ell\in\mathcal{S}(\mathrm{M}_{3})\mid\ell(0\oplus 1)>0\}$ contains
$\omega_{\lambda}:=r_{\mathrm{M}_{3}}\big{[}(1-\lambda)\ket{+}\!\!\bra{+}\oplus\lambda\big{]}\,,\quad\lambda\in(0,1]\,.$
So $\mathcal{O}|_{\mathcal{R}}:=\{\ell|_{\mathcal{R}}:\ell\in\mathcal{O}\}$ contains $\omega_{\lambda}|_{\mathcal{R}}$ but none of the restrictions $\ell_{\theta}|_{\mathcal{R}}$ of
$\ell_{\theta}:=r_{\mathrm{M}_{3}}\big{[}\frac{1}{2}\big{(}\mathds{1}_{2}+\cos(\theta)X+\sin(\theta)Z\big{)}\oplus 0\big{]}\,,\quad\theta\in(0,2\pi)\,.$
Specifically, $\ell_{\theta}$ has the value $\ell_{\theta}(A_{\theta})=1$ at $A_{\theta}:=\cos(\theta)(X\oplus 1)+\sin(\theta)Z\oplus 0$ and $\ell(A_{\theta})\leq(\cos(\theta)-1)\ell(0\oplus 1)+1$ holds for every $\ell\in\mathcal{S}(\mathrm{M}_{3})$. Hence
$\ell|_{\mathcal{R}}(A_{\theta})-\ell_{\theta}|_{\mathcal{R}}(A_{\theta})\leq(\cos(\theta)-1)\ell(0\oplus 1)<0\,,\quad\ell\in\mathcal{O}\,.$
This shows that $\mathcal{O}|_{\mathcal{R}}$ is not a neighborhood of $\omega_{\lambda}|_{\mathcal{R}}$ as $\lim_{\theta\to 0}\ell_{\theta}|_{\mathcal{R}}=\omega_{\lambda}|_{\mathcal{R}}$. In conclusion, $\mathcal{S}(\mathrm{M}_{3})\to\mathcal{S}(\mathcal{R})$, $\ell\mapsto\ell|_{\mathcal{R}}$ is not open at $\omega_{\lambda}$ for any $\lambda\in(0,1]$.
Asking where $\mathcal{S}(\mathrm{M}_{n})\to\mathcal{S}(\mathcal{R})$, $\ell\mapsto\ell|_{\mathcal{R}}$ is open is the same, by Coro. 2.7 b), as inquiring at which density matrices the orthogonal projection
$\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R})$ (1.1)
is open. The map (1.1) is defined in (2.2) and (2.3) below. For now suffice it to say that it is a restriction of the orthogonal projection of $\mathrm{M}_{n}$ onto $\mathcal{R}$.
Corey et al. [17] and Leake et al. [29, 30] first studied a problem of numerical ranges closely related to (1.1). The problem (1.1) itself was studied by Weis [53, 54] and Rodman et al. [46] when $\mathcal{D}(\mathcal{R})$ is replaced with the affinely isomorphic joint numerical range. Numerical ranges are the topic of Sec. 4.
Theorem 1.2.
If $\mathcal{A}$ is a *-subalgebra of $\mathrm{M}_{n}$, then the orthogonal projection $\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{A})$ is open.
Thm. 1.2 can simplify the problem (1.1), and related continuity problems, if $\mathcal{R}$ is included in a *-subalgebra of $\mathrm{M}_{n}$ smaller than $\mathrm{M}_{n}$ as we show in Sec. 8. Examples from quantum information theory are presented in Sec. 9.
Thm. 1.2 is proved in Sec. 6. The main ideas are that $\mathcal{A}$ is a direct sum of full matrix algebras and that $\mathcal{D}(\mathrm{M}_{n})$ is stable and highly symmetric. Thereby, a convex set $K$ is stable if the midpoint map
$K\times K\to K\,,\quad(x,y)\mapsto\frac{1}{2}(x+y)$
is open. A convex set is always (except Rem. 1.3) understood to be included in a Euclidean space. Debs [20] proved the stability of $\mathcal{D}(\mathcal{A})$ for $\mathcal{A}=\mathrm{M}_{n}$, Papadopoulou [41] achieved it for *-subalgebras $\mathcal{A}$ of $\mathrm{M}_{n}$, and we do it for real *-subalgebras $\mathcal{A}$ of $\mathrm{M}_{n}$ in Sec. 5 using Clausing’s work on retractions [15]. The analogue of Thm. 1.2 is established for the algebra $\mathcal{A}=\mathrm{M}_{n}(\mathbb{R})$ of real $n\times n$ matrices in Sec. 7. The Secs. 2 and 3 collect preliminaries.
Remark 1.3.
  1. a)
    Vesterstrøm [52] proved that the restriction of states to the center of a von Neumann algebra is open. Thm. 1.2 is a noncommutative analogue in the finite-dimensional setting.
  2. b)
    Stability is a meaningful concept in optimal control [41, 58] and quantum information theory [48] because of the following “CE-property”. A compact convex set $K$ is stable if and only if for every continuous function $f:K\to\mathbb{R}$, the envelope $f^{\vee}(x):=\sup\{h(x):h\leq f\}$, $x\in K$ is continuous, see [52, 39] and [48]. Here, the supremum is taken over all continuous affine functions $h:K\to\mathbb{R}$ whose graphs lie below the graph of $f$.
2. States and density matrices
The aim of this section is to translate openness questions from states to density matrices. In the sequel, we refer to matrix systems and *-subalgebras synonymously as complex matrix systems and complex *-subalgebras, respectively. In addition, we introduce real *-subalgebras as they have a greater variety of state spaces than the complex ones. See, however, Rem. 2.4.
A real matrix system on $\mathbb{C}^{n}$ is a real linear subspace $\mathcal{R}$ of $\mathrm{M}_{n}$ that is self-adjoint and contains $\mathds{1}_{n}$. Let $\mathcal{R}$ denote a real matrix system on $\mathbb{C}^{n}$. We endow $\mathcal{R}$ with the Euclidean scalar product
×$\mathcal{R}\times\mathcal{R}\to\mathbb{R}\,,\quad(A,B)\mapsto\mathrm{Re}\langle A,B\rangle\,,$ (2.1)
where $\langle A,B\rangle=\operatorname{tr}A^{\ast}B$ and $\mathrm{Re}(a+\mathrm{i}b)=a$ is the real part of a complex number, $A,B\in\mathrm{M}_{n}$, $a,b\in\mathbb{R}$. The positive cone, space of hermitian matrices, dual cone, and space of density matrices are defined verbatim to their respective complex counterparts defined in Sec. 1, and are denoted by
$\mathcal{C}(\mathcal{R}),\quad\mathcal{H}(\mathcal{R}),\quad\mathcal{C}(\mathcal{R})^{\vee},\quad\text{and}\quad\mathcal{D}(\mathcal{R})\,.$
We call $\mathcal{R}$ a real *-subalgebra of $\mathrm{M}_{n}$ if $\mathcal{R}$ is closed under matrix multiplication.
Generalizing a definition from real algebras [32, Sec. 4.5], we define the real state space of $\mathcal{R}$ as
$\mathcal{S}_{\mathbb{R}}(\mathcal{R}):=\{\ell\in\mathcal{R}_{\mathbb{R},0}^{\ast}\mid\forall A\in\mathcal{C}(\mathcal{R}):\ell(A)\geq 0,\ell(\mathds{1}_{n})=1\}\,,$
where
$\mathcal{R}_{\mathbb{R},0}^{\ast}:=\{\ell:\mathcal{R}\to\mathbb{R}\mid\mbox{$\ell$ is $\mathbb{R}$-linear},\forall A\in\mathcal{H}^{-}(\mathcal{R}):\ell(A)=0\}$
is the space of real functionals vanishing on the skew-hermitian matrices
():=$\mathcal{H}^{-}(\mathcal{R}):=\{A\in\mathcal{R}:A^{\ast}=-A\}\,.$
Lemma 2.1.
If $\mathcal{R}$ is a real matrix system on $\mathbb{C}^{n}$, then the map
$r_{\mathbb{R},\mathcal{R}}:\mathcal{D}(\mathcal{R})\to\mathcal{S}_{\mathbb{R}}(\mathcal{R}),\quad\rho\mapsto\mathrm{Re}\langle\rho,\,\cdot\,\rangle$
is a real affine isomorphism between compact convex sets.
Proof.
As $\mathcal{R}$ is the orthogonal direct sum $\mathcal{R}=\mathcal{H}(\mathcal{R})\oplus\mathcal{H}^{-}(\mathcal{R})$, the real linear isomorphism [23, Sec. 67]
$r_{\mathbb{R},\mathcal{R}}:\mathcal{R}\to\{\ell:\mathcal{R}\to\mathbb{R}\mid\mbox{$\ell$ is $\mathbb{R}$-linear}\}\,,\quad A\mapsto\mathrm{Re}\langle A,\,\cdot\,\rangle$
restricts to the real linear isomorphism $\mathcal{H}(\mathcal{R})\to\mathcal{R}_{\mathbb{R},0}^{\ast}$. Restricting $r_{\mathbb{R},\mathcal{R}}$ further, we obtain an injective map whose domain is $\mathcal{D}(\mathcal{R})$. So, it suffices to prove $r_{\mathbb{R},\mathcal{R}}(\rho)\in\mathcal{S}_{\mathbb{R}}(\mathcal{R})$ for $\rho\in\mathcal{D}(\mathcal{R})$, and that $r_{\mathbb{R},\mathcal{R}}:\mathcal{D}(\mathcal{R})\to\mathcal{S}_{\mathbb{R}}(\mathcal{R})$ is surjective. Both assertions are straightforward to verify. The convex set $\mathcal{D}(\mathcal{R})$ is compact, see [44, Lemma 3.3], because $\mathds{1}_{n}$ lies in the interior of $\mathcal{C}(\mathcal{R})$ in the topology of $\mathcal{H}(\mathcal{R})$. The convex set $\mathcal{S}_{\mathbb{R}}(\mathcal{R})$ is compact as it is the image of a compact set under a continuous map. ∎
Returning to complex functionals, we consider the real vector space
$\mathcal{R}_{\mathrm{her}}^{\ast}:=\{\ell:\mathcal{R}\to\mathbb{C}\mid\mbox{$\ell$ is $\mathbb{C}$-linear},\forall A\in\mathcal{H}(\mathcal{R}):\ell(A)\in\mathbb{R}\}$
of complex functionals taking real values on hermitian matrices.
Lemma 2.2.
If $\mathcal{R}$ is a complex matrix system on $\mathbb{C}^{n}$, then the map
$r_{\mathcal{R}}:\mathcal{D}(\mathcal{R})\to\mathcal{S}(\mathcal{R})\,,\quad\rho\mapsto\langle\rho,\,\cdot\,\rangle$
is a real affine isomorphism between compact convex sets.
Proof.
The complex antilinear isomorphism $r_{\mathcal{R}}:\mathcal{R}\to\mathcal{R}^{\ast}$, $A\mapsto\langle A,\,\cdot\,\rangle$ restricts to a real linear isomorphism $\mathcal{H}(\mathcal{R})\to\mathcal{R}_{\mathrm{her}}^{\ast}$. Furthermore,
$\alpha:\mathcal{R}_{\mathbb{R},0}^{\ast}\to\mathcal{R}_{\mathrm{her}}^{\ast}\,,\quad\alpha(\ell)[A+\mathrm{i}B]:=\ell(A)+\mathrm{i}\ell(B)\,,\quad A,B\in\mathcal{H}(\mathcal{R})\,,$
is a real linear isomorphism, whose inverse is given by $\ell\mapsto\mathrm{Re}\circ\ell$. So, the following diagram commutes. (Note that two arrows in opposite directions denote a bijection.)
${\mathcal{H}(\mathcal{R})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r_{\mathcal{R}}}$$\scriptstyle{r_{\mathbb{R},\mathcal{R}}}$${\mathcal{R}_{\mathbb{R},0}^{\ast}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha}$${\mathcal{R}_{\mathrm{her}}^{\ast}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{Re}}$
Moreover, if $\ell_{1}\in\mathcal{R}_{\mathbb{R},0}^{\ast}$ and $\ell_{2}\in\mathcal{R}_{\mathrm{her}}^{\ast}$ satisfy $\ell_{2}=\alpha(\ell_{1})$, then $\ell_{1}\in\mathcal{S}_{\mathbb{R}}(\mathcal{R})$ holds if and only if $\ell_{2}\in\mathcal{S}(\mathcal{R})$. Hence, the claim follows from Lemma 2.1. ∎
Example 2.3.
Real *-subalgebras of $\mathrm{M}_{n}$ have a richer class of state spaces than the complex ones. The Bloch ball [7, Sec. 5.2]
$\mathcal{D}(\mathrm{M}_{2})=\big{\{}\frac{1}{2}(\mathds{1}_{2}+c_{X}X+c_{Y}Y+c_{Z}Z)\colon c_{X},c_{Y},c_{Z}\in\mathbb{R},c_{X}^{2}+c_{Y}^{2}+c_{Z}^{2}\leq 1\big{\}}$
is a three-dimensional Euclidean ball of radius $1/\sqrt{2}$. The set of density matrices of the algebra $\mathrm{M}_{2}(\mathbb{R})$ of real $2\times 2$ matrices is the great disk
$\mathcal{D}(\mathrm{M}_{2}(\mathbb{R}))=\big{\{}\frac{1}{2}(\mathds{1}_{2}+c_{X}X+c_{Z}Z)\colon c_{X},c_{Z}\in\mathbb{R},c_{X}^{2}+c_{Z}^{2}\leq 1\big{\}}$
of the Bloch ball $\mathcal{D}(\mathrm{M}_{2})$. There is no complex *-subalgebra of $\mathrm{M}_{n}$ in any dimension $n$ whose state space is a disk.
Remark 2.4.
Real and complex matrix systems have the same families of state spaces. If $\mathcal{R}$ is a real or complex matrix system on $\mathbb{C}^{n}$, then the real matrix system of hermitian matrices $\mathcal{H}(\mathcal{R})$ has the same set of hermitian matrices and hence the same set of density matrices as $\mathcal{R}$. Conversely, any real matrix system $\mathcal{R}$ included in $\mathcal{H}(\mathrm{M}_{n})$ has the same same set of hermitian matrices and the same set of density matrices as the complex matrix system $\mathcal{R}\oplus\mathrm{i}\mathcal{R}$. The set of density matrices of a real or complex matrix system $\mathcal{R}$ is affinely isomorphic to the state space of $\mathcal{R}$ by Lemma 2.1 or 2.2, respectively.
Let $\mathcal{R}_{1},\mathcal{R}_{2}$ be real matrix systems on $\mathbb{C}^{n}$ and let $\mathcal{R}_{2}\subset\mathcal{R}_{1}$. We abbreviate $\mathcal{H}_{i}:=\mathcal{H}(\mathcal{R}_{i})$, $\mathcal{C}_{i}:=\mathcal{C}(\mathcal{R}_{i})$, $\mathcal{C}_{i}^{\vee}:=\mathcal{C}^{\vee}(\mathcal{R}_{i})$, $\mathcal{D}_{i}:=\mathcal{D}(\mathcal{R}_{i})$, $\mathcal{S}_{\mathbb{R},i}:=\mathcal{S}_{\mathbb{R}}(\mathcal{R}_{i})$, $\mathcal{S}_{i}:=\mathcal{S}(\mathcal{R}_{i})$, $r_{\mathbb{R},i}:=r_{\mathbb{R},\mathcal{R}_{i}}$, and $r_{i}:=r_{\mathcal{R}_{i}}$, $i=1,2$. Usually, the orthogonal projection of $\mathcal{R}_{1}$ onto $\mathcal{R}_{2}$ is the idempotent self-adjoint linear map $\mathcal{R}_{1}\to\mathcal{R}_{1}$ whose range is $\mathcal{R}_{2}$. Reducing the codomain to the range, we get a map
$\pi:\mathcal{R}_{1}\to\mathcal{R}_{2}\,,$ (2.2)
whose value at $A\in\mathcal{R}_{1}$ is specified by the equations $\mathrm{Re}\langle A-\pi(A),B\rangle=0$ for all $B\in\mathcal{R}_{2}$. We refer to $\pi$ as the orthogonal projection of $\mathcal{R}_{1}$ onto $\mathcal{R}_{2}$ in this paper. The notation $\mathcal{R}_{1}\to\mathcal{R}_{2}$ conveying domain and range is useful especially in Sec. 6. The adjoint of $\pi$ is the embedding $\mathcal{R}_{2}\to\mathcal{R}_{1}$, $A\mapsto A$. If $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ are complex matrix systems, then the Frobenius inner product induces the same orthogonal projection as the Euclidean scalar product (2.1).
As $\mathcal{R}_{i}$ is the orthogonal direct sum of the spaces of its hermitian and skew-hermitian matrices, $i=1,2$, the map (2.2) restricts to $\pi:\mathcal{H}_{1}\to\mathcal{H}_{2}$, which we denote (aware of the notational imprecision) by the same symbol $\pi$. The value of $\pi$ at $A\in\mathcal{H}_{1}$ is specified by $\langle A-\pi(A),B\rangle=0$ for all $B\in\mathcal{H}_{2}$.
A cone in a Euclidean space $(E,\langle\!\langle\cdot,\cdot\rangle\!\rangle)$ is a subset $C$ of $E$ that is closed under multiplication with positive scalars. A base of a cone $C$ is a subset $B$ of $C$ such that $0\not\in\operatorname{aff}B$ and such that for all nonzero $x\in C$ there exist $y\in B$ and $s>0$ such that $x=sy$ holds. Note that we have $B=C\cap\operatorname{aff}B$ for every base $B$ of a cone $C$. The set $M^{\vee}:=\{x\in E\mid\forall y\in M\colon\langle\!\langle x,y\rangle\!\rangle\geq 0\}$ is a closed convex cone for every subset $M\subset E$, called the dual cone to $M$. Regarding duality of convex cones, we refer to [45, Sec. 14].
Lemma 2.5.
Let $\mathcal{R}_{1},\mathcal{R}_{2}$ be real matrix systems on $\mathbb{C}^{n}$ such that $\mathcal{R}_{2}\subset\mathcal{R}_{1}$ and let $\pi:\mathcal{H}_{1}\to\mathcal{H}_{2}$ denote the orthogonal projection. Then $\mathcal{C}_{2}^{\vee}=\pi(\mathcal{C}_{1}^{\vee})$ and $\mathcal{D}_{2}=\pi(\mathcal{D}_{1})$ holds.
Proof.
This lemma and its proof are similar to [44, Prop. 5.2]. The inclusions $\pi(\mathcal{C}_{1}^{\vee})\subset\mathcal{C}_{2}^{\vee}$ and $\mathcal{C}_{2}\supset(\pi(\mathcal{C}_{1}^{\vee}))^{\vee}$ are straightforward to verify and imply
$\pi(\mathcal{C}_{1}^{\vee})\subset\mathcal{C}_{2}^{\vee}\subset[\pi(\mathcal{C}_{1}^{\vee})]^{\vee}{}^{\vee}\,.$
Since $[\pi(\mathcal{C}_{1}^{\vee})]^{\vee}{}^{\vee}$ is the closure of the convex cone $\pi(\mathcal{C}_{1}^{\vee})$, it suffices to show that $\pi(\mathcal{C}_{1}^{\vee})$ is closed. By [44, Lemma 3.1], this would follow if $\pi(\mathcal{D}_{1})$ was a compact base of $\pi(\mathcal{C}_{1}^{\vee})$. As $\mathds{1}_{n}$ lies in the interior of $\mathcal{C}_{1}$ in the topology of $\mathcal{H}_{1}$, we know that $\mathcal{D}_{1}$ is a compact base of $\mathcal{C}_{1}^{\vee}$, see [44, Lemma 3.3]. Hence $\pi(\mathcal{D}_{1})$ is a compact base of $\pi(\mathcal{C}_{1}^{\vee})$ provided we establish that $0\not\in\operatorname{aff}\pi(\mathcal{D}_{1})$. But this is clear from $\pi$ being trace-preserving:
$\operatorname{tr}(\pi(A))=\langle\mathds{1}_{n},\pi(A)\rangle=\langle\mathds{1}_{n},A\rangle=\operatorname{tr}(A)\,,\quad A\in\mathcal{H}_{1}\,.$
This completes the proof of $\mathcal{C}_{2}^{\vee}=\pi(\mathcal{C}_{1}^{\vee})$. The identity $\mathcal{C}_{2}^{\vee}=\pi(\mathcal{C}_{1}^{\vee})$ and the fact that $\pi$ is trace-preserving imply $\mathcal{D}_{2}=\pi(\mathcal{D}_{1})$. ∎
Lemma 2.5 shows that the orthogonal projection $\pi:\mathcal{R}_{1}\to\mathcal{R}_{2}$ restricts to the map
$\pi:\mathcal{D}_{1}\to\mathcal{D}_{2}\,,$ (2.3)
which we call the orthogonal projection of $\mathcal{D}_{1}$ onto $\mathcal{D}_{2}$. To avoid any possible confusion, the orthogonal projections $\mathcal{R}_{1}\to\mathcal{R}_{2}$ and $\mathcal{H}_{1}\to\mathcal{H}_{2}$ are written without function labels from here on (Lemma 3.7 is an exception).
Proposition 2.6.
Let $\mathcal{R}_{1},\mathcal{R}_{2}$ be real matrix systems on $\mathbb{C}^{n}$ such that $\mathcal{R}_{2}\subset\mathcal{R}_{1}$ and let $\pi:\mathcal{D}(\mathcal{R}_{1})\to\mathcal{D}(\mathcal{R}_{2})$ denote the orthogonal projection.
  1. a)
    The diagram a) below commutes. Both $\pi:\mathcal{D}_{1}\to\mathcal{D}_{2}$ and $\mathcal{S}_{\mathbb{R},1}\to\mathcal{S}_{\mathbb{R},2}$, $\ell\mapsto\ell|_{\mathcal{R}_{2}}$ are surjective maps.
  2. b)
    If $\mathcal{R}_{1},\mathcal{R}_{2}$ are complex matrix systems on $\mathbb{C}^{n}$, then the diagram b) below commutes and the maps $\pi:\mathcal{D}_{1}\to\mathcal{D}_{2}$ and $\mathcal{S}_{1}\to\mathcal{S}_{2}$, $\ell\mapsto\ell|_{\mathcal{R}_{2}}$ are onto.
a)   ${\mathcal{D}_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r_{\mathbb{R},1}}$$\scriptstyle{\pi}$${\mathcal{S}_{\mathbb{R},1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\ell\mapsto\ell|_{\mathcal{R}_{2}}}$${\mathcal{D}_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r_{\mathbb{R},2}}$${\mathcal{S}_{\mathbb{R},2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$            b)   ${\mathcal{D}_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r_{1}}$$\scriptstyle{\pi}$${\mathcal{S}_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\ell\mapsto\ell|_{\mathcal{R}_{2}}}$${\mathcal{D}_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r_{2}}$${\mathcal{S}_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
Proof.
a) The horizontal arrows of diagram a) and b) are obtained in Lemma 2.1 and Lemma 2.2, respectively, and $\pi:\mathcal{D}_{1}\to\mathcal{D}_{2}$ is onto by Lemma 2.5. The diagram a) commutes because for all $\rho\in\mathcal{D}_{1}$ and $A\in\mathcal{R}_{2}$ we have
$[r_{\mathbb{R},2}\circ\pi(\rho)](A)=\mathrm{Re}\langle\pi(\rho),A\rangle=\mathrm{Re}\langle\rho,A\rangle=[r_{\mathbb{R},1}(\rho)](A)=r_{\mathbb{R},1}(\rho)|_{\mathcal{R}_{2}}(A)\,.$
Therefore and since $\pi:\mathcal{D}_{1}\to\mathcal{D}_{2}$ is surjective, the map $\mathcal{S}_{\mathbb{R},1}\to\mathcal{S}_{\mathbb{R},2}$ is surjective as well. The proof of b) is similar. ∎
Corollary 2.7.
Let $\mathcal{R}_{1},\mathcal{R}_{2}$ be real matrix systems on $\mathbb{C}^{n}$ such that $\mathcal{R}_{2}\subset\mathcal{R}_{1}$, let $\pi:\mathcal{D}(\mathcal{R}_{1})\to\mathcal{D}(\mathcal{R}_{2})$ denote the orthogonal projection, and let $\rho\in\mathcal{D}_{1}$.
  1. a)
    The map $\pi:\mathcal{D}_{1}\to\mathcal{D}_{2}$ is open at $\rho$ if and only if $\mathcal{S}_{\mathbb{R},1}\to\mathcal{S}_{\mathbb{R},2}$, $\ell\mapsto\ell|_{\mathcal{R}_{2}}$ is open at $r_{\mathbb{R},1}(\rho)$.
  2. b)
    If $\mathcal{R}_{1},\mathcal{R}_{2}$ are complex matrix systems on $\mathbb{C}^{n}$, then $\pi:\mathcal{D}_{1}\to\mathcal{D}_{2}$ is open at $\rho$ if and only if $\mathcal{S}_{1}\to\mathcal{S}_{2}$, $\ell\mapsto\ell|_{\mathcal{R}_{2}}$ is open at $r_{1}(\rho)$.
Proof.
This follows directly from Prop. 2.6. ∎
Remark 2.8.
Let $\mathcal{R}_{1},\mathcal{R}_{2}$ be real matrix systems on $\mathbb{C}^{n}$ and let $\mathcal{R}_{2}\subset\mathcal{R}_{1}$. Then $\mathcal{D}_{2}\supset\mathcal{D}_{1}\cap\mathcal{R}_{2}$ holds, but the converse inclusion is wrong in general, as Ex. 2.9 shows. However, if $\mathcal{R}_{2}$ is a real *-subalgebra of $\mathrm{M}_{n}$, then we have
$\mathcal{D}_{2}=\mathcal{D}_{1}\cap\mathcal{R}_{2}\,.$ (2.4)
Indeed, $\mathcal{C}_{2}^{\vee}=\mathcal{C}_{2}$ holds [28, Thm. III.2.1] as the space of hermitian matrices $\mathcal{H}_{2}$ is a Euclidean Jordan algebra with Jordan product $A\circ B=\frac{1}{2}(AB+BA)$ and inner product $(A,B)\mapsto\mathrm{Re}\langle A,B\rangle$, $A,B\in\mathcal{H}_{2}$. So, $\mathcal{C}_{2}^{\vee}=\mathcal{C}_{2}\subset\mathcal{C}_{1}\subset\mathcal{C}_{1}^{\vee}$ proves $\mathcal{D}_{2}\subset\mathcal{D}_{1}$, which implies (2.4). See also the Notes to Chapter 6 in [2].
Example 2.9.
Despite $\mathcal{R}_{2}\subset\mathcal{R}_{1}\subset\mathrm{M}_{3}$, the inclusions $\mathcal{D}_{2}\subset\mathcal{D}_{1}\subset\mathcal{D}(\mathrm{M}_{3})$ fail if $\mathcal{R}_{2}:=\operatorname{span}_{\mathbb{C}}(\mathds{1}_{3},Z\oplus 0)$ and $\mathcal{R}_{1}:=\operatorname{span}_{\mathbb{C}}(\mathds{1}_{3},X\oplus 1,Z\oplus 0)$. We have
$A_{\lambda}\in\mathcal{D}_{2}\Leftrightarrow|\lambda|\leq\frac{3}{2}\,,\quad A_{\lambda}\in\mathcal{D}_{1}\Leftrightarrow|\lambda|\leq\sqrt{2}\,,\quad A_{\lambda}\in\mathcal{D}(\mathrm{M}_{3})\Leftrightarrow|\lambda|\leq 1$
for $A_{\lambda}:=(\mathds{1}_{3}+\lambda\,Z\oplus 0)/3\in\mathcal{R}_{2}$, $\lambda\in\mathbb{R}$. The second equivalence is obtained by minimizing $\langle A_{\lambda},A\rangle$ for fixed $\lambda$ over $A\in\mathcal{C}_{1}$, that is, by minimizing $\langle A_{\lambda},\mathds{1}_{3}+c_{1}(X\oplus 1)+c_{2}Z\oplus 0\rangle=1+\frac{1}{3}c_{1}+\frac{2}{3}\lambda c_{2}$ on the unit disk of $\mathbb{R}^{2}$.
3. Direct convex sums
This section addresses affinely independent convex sets, their convex hulls, and maps defined thereon. Let
$\Delta_{m}:=\{(s_{1},\dots,s_{m})\in\mathbb{R}^{m}\mid\forall i:s_{i}\geq 0,s_{1}+\dots+s_{m}=1\}$
denote the probability simplex, and
$\Delta_{m}(\epsilon,s_{1},\dots,s_{m}):=\Delta_{m}\cap\bigoplus_{i=1}^{m}(s_{i}-\epsilon,s_{i}+\epsilon)$
the open hypercube of edge length $2\epsilon$ centered at $(s_{1},\dots,s_{m})\in\Delta_{m}$.
A family of convex subsets $K_{1},\dots,K_{m}$ of a Euclidean space is affinely independent if every point in their convex hull can be expressed by a unique convex combination $s_{1}x_{1}+\dots+s_{m}x_{m}$. This means that $(s_{1},\dots,s_{m})\in\Delta_{m}$ is unique and $x_{i}\in K_{i}$ is unique for all $i$ for which if $s_{i}>0$, $i=1,\dots,m$. The direct convex sum [1] of a family $K_{1},\dots,K_{m}$ of affinely independent convex sets is defined as their convex hull
$K_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}K_{m}:=\operatorname{conv}(K_{1}\cup\dots\cup K_{m})\,.$
If $K_{1},\dots,K_{m}$ is a family of affinely independent compact convex sets, then their direct convex sum is compact [45, Thm. 17.2]. A compactness argument allows us to describe a base of open neighborhoods.
Lemma 3.1.
Let $K_{1},\dots,K_{m}$ be affinely independent compact convex subsets of a Euclidean space, and let $x_{i}\in K_{i}$, $i=1,\dots,m$, and $(s_{1},\dots,s_{m})\in\Delta_{m}$. Let $I:=\{i\in\{1,\ldots,m\}\mid s_{i}>0\}$, $\delta:=\min_{i\in I}s_{i}$, and define
$O_{I}(\epsilon,(A_{i})_{i\in I}):=\{t_{1}y_{1}+\dots+t_{m}y_{m}\mid\hskip 2.0&(t_{1},\dots,t_{m})\in\Delta_{m}(\epsilon,s_{1},\dots,s_{m}),\\ \hskip 2.0&\mbox{$\forall i\colon y_{i}\in K_{i}$ and ($i\in I\Rightarrow y_{i}\in A_{i}$)}\}\end{array}$
for every $\epsilon\in(0,\delta]$ and open set $A_{i}$ in the relative topology of $K_{i}$ containing $x_{i}$, $i=1,\dots,m$. Then the family $\{O_{I}(\epsilon,(A_{i})_{i\in I})\}$ is a local base of open neighborhoods at $s_{1}x_{1}+\dots+s_{m}x_{m}$ in the relative topology of $K_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}K_{m}$.
Proof.
As $K:=K_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}K_{m}$ is a metric space, it suffices to show that $O:=O_{I}(\epsilon,(A_{i})_{i\in I})$ is open and there are arbitrary small sets of this form.
Let $B_{i}:=A_{i}$ if $i\in I$ and $B_{i}:=K_{i}$ if $i\not\in I$, $i=1,\dots,m$. Then the set
$U:=\Delta_{m}(\epsilon,s_{1},\dots,s_{m})\oplus\bigoplus_{i}B_{i}$
is open in the relative topology of the compact set $\widetilde{K}:=\Delta_{m}\oplus\bigoplus_{i}K_{i}$. The complement
$U^{\complement}:=\widetilde{K}\setminus U=\underbrace{\textstyle\Delta_{m}(\epsilon,s_{1},\dots,s_{m})^{\complement}\oplus\bigoplus_{i}K_{i}}_{C:=}\cup\bigcup_{j\in I}\underbrace{\textstyle\Delta_{m}\oplus(\bigoplus_{i\neq j}K_{i})\oplus A_{j}^{\complement}}_{C_{j}:=}$
is compact. Hence, its image under the continuous surjective map
$\beta:\widetilde{K}\to K\,,\quad((t_{i})_{i=1}^{m},(y_{i})_{i=1}^{m})\mapsto\sum_{i=1}^{m}t_{i}y_{i}$
is compact. We prove that $O=\beta(U)$ is open in $K$ by showing that $\beta(U)$ is disjoint from $\beta(U^{\complement})$. Let
$u:=(t_{1},\dots,t_{m},y_{1},\dots,y_{m})\in U\,.$
Then $\beta(u)\not\in\beta(C)$ follows as the vector $(t_{i})_{i=1}^{m}$ is uniquely determined by $\beta(u)$. For every $j\in I$ we have $s_{j}>0$, hence $t_{j}>0$ by the definition of $O$. Thus, $y_{j}$ is uniquely determined by $\beta(u)$, which proves $\beta(u)\not\in\beta(C_{j})$.
Let $|M|:=\sup_{x,y\in M}|y-x|$ denote the diameter of a set $M$. The distance of $x:=s_{1}x_{1}+\dots+s_{m}x_{m}$ from a point $t_{1}y_{1}+\dots+t_{m}y_{m}$ in $O$ is bounded by
$\sum_{i}\left(|s_{i}-t_{i}||x_{i}|+t_{i}|x_{i}-y_{i}|\right)\leq\epsilon\sum_{i}|x_{i}|+\sum_{i\in I}t_{i}|A_{i}|+\epsilon\sum_{i\not\in I}|K_{i}|\,.$
Choosing open sets $(A_{i})_{i\in I}$ with diameters at most $\epsilon$, we obtain
$|O|\leq 2\sup_{y\in O}|x-y|\leq 2\epsilon\left(\sum_{i}|x_{i}|+1+\sum_{i\not\in I}|K_{i}|\right)\,,$
which completes the proof, as the compact sets $K_{i}$ have finite diameters and as $\epsilon$ can be chosen arbitrarily small. ∎
Let $E_{1},\dots,E_{m}$ be Euclidean spaces. We consider $E_{i}$ as a subspace of the direct sum $\bigoplus_{j=1}^{m}E_{j}$ via the embedding
$E_{i}\to\bigoplus_{j=1}^{m}E_{j}\,,\quad x\mapsto(\underbrace{0,\dots,0}_{\makebox[0.0pt]{\scriptsize$i-1$ zeros}},x,\underbrace{0,\dots,0}_{\makebox[0.0pt]{\scriptsize$m-i$ zeros}})\,,\quad i=1,\dots,m\,.$
Proposition 3.2.
Let $K_{i}\subset E_{i}$ and $L_{i}\subset F_{i}$ be compact convex subsets of Euclidean spaces $E_{i}$ and $F_{i}$, $i=1,\dots,m$, such that $K_{1},\dots,K_{m}$ are affinely independent in $\bigoplus_{i=1}^{m}E_{i}$ and $L_{1},\dots,L_{m}$ are affinely independent in $\bigoplus_{i=1}^{m}F_{i}$. If $f_{i}:K_{i}\to L_{i}$ is a map, $i=1,\dots,m$, then a map
$f_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}f_{m}:K_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}K_{m}\to L_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}L_{m}$ (3.1)
is well defined by
$s_{1}x_{1}+\dots+s_{m}x_{m}\mapsto s_{1}f_{1}(x_{1})+\dots+s_{m}f_{m}(x_{m})\,.$
If $f_{i}$ is open and surjective for $i=1,\dots,m$, then $f_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}f_{m}$ is open.
Proof.
The map is well defined as the sets $K_{1},\dots,K_{m}$ are affinely independent and the sets $L_{1},\dots,L_{m}$ are affinely independent. Regarding the openness of $f_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}f_{m}$, it suffices to show that the images of the members of a base of the relative topology of $K_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}K_{m}$ are open. Using the base of Lemma 3.1, we have
$f_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}f_{m}[O_{I}(\epsilon,(A_{i})_{i\in I})]=O_{I}[\epsilon,(f(A_{i}))_{i\in I}]\,,$
which is open in $L_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}L_{m}$. The required identity of $f_{i}(K_{i})=L_{i}$ for every $i\not\in I$ is a consequence of $f_{i}$ being surjective. ∎
We call the map (3.1) the direct convex sum of the maps $f_{1},\dots,f_{m}$. Next, we recall a sufficient condition for the affine independence of convex sets [16].
Remark 3.3.
Let $K_{i}$ be a convex subset of a Euclidean space $E_{i}$ such that $0\not\in\operatorname{aff}K_{i}$ holds for $i=1,\dots,m$. Then $K_{1},\dots,K_{m}$ are affinely independent in the direct sum $\bigoplus_{i=1}^{m}E_{i}$.
Returning to matrix systems, we consider a real matrix system $\mathcal{R}_{i}$ on $\mathbb{C}^{n_{i}}$, $i=1,\dots,m$. The direct sum $\mathcal{R}:=\bigoplus_{i=1}^{m}\mathcal{R}_{i}$ is a real matrix system on $\mathbb{C}^{n_{1}+\dots+n_{m}}$. The Frobenius inner product of $(A_{i})_{i=1}^{m},(B_{i})_{i=1}^{m}\in\mathcal{R}$ is $\langle(A_{i})_{i=1}^{m},(B_{i})_{i=1}^{m}\rangle=\sum_{i=1}^{m}\langle A_{i},B_{i}\rangle$.
Lemma 3.4.
Let $\mathcal{R}_{i}$ be a real matrix system on $\mathbb{C}^{n_{i}}$, $i=1,\ldots,m$. Then
$\mathcal{C}(\mathcal{R}_{1}\oplus\dots\oplus\mathcal{R}_{m})$ $=\mathcal{C}(\mathcal{R}_{1})\oplus\dots\oplus\mathcal{C}(\mathcal{R}_{m})\,,$
$\mathcal{C}^{\vee}(\mathcal{R}_{1}\oplus\dots\oplus\mathcal{R}_{m})$ $=\mathcal{C}^{\vee}(\mathcal{R}_{1})\oplus\dots\oplus\mathcal{C}^{\vee}(\mathcal{R}_{m})\,,$
$\text{and}\quad\mathcal{D}(\mathcal{R}_{1}\oplus\dots\oplus\mathcal{R}_{m})$ $=\mathcal{D}(\mathcal{R}_{1})\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}\mathcal{D}(\mathcal{R}_{m})\,.$
Proof.
The first identity is clear. The second one follows by induction from $m=2$, a case that is easy to verify. By Rem. 3.3 and because $\mathcal{D}(\mathcal{R}_{i})$ is included in the hyperplane of trace-one matrices $i=1,\dots,m$, the sets $\mathcal{D}(\mathcal{R}_{1}),\dots,\mathcal{D}(\mathcal{R}_{m})$ are affinely independent in the direct sum $\bigoplus_{i=1}^{m}\mathcal{H}(\mathcal{R}_{i})$. The third identity follows from the second one by enforcing the trace to be one. ∎
Example 3.5.
Let $\mathcal{R}:=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{3},X\oplus 1,Z\oplus 0)$.
  1. a)
    The set of density matrices $\mathcal{D}(\mathrm{M}_{2}\oplus\mathrm{M}_{1})$ of the *-subalgebra $\mathrm{M}_{2}\oplus\mathrm{M}_{1}$ of $\mathrm{M}_{3}$ is a symmetric cone that is the direct convex sum of the Bloch ball $\mathcal{D}(\mathrm{M}_{2})$ and the singleton $\mathcal{D}(\mathrm{M}_{1})=\{1\}$, see Lemma 3.4 and Ex. 2.3. The closed segment
    $\mathcal{G}:=\big{[}\ket{+}\!\!\bra{+}\oplus 0,0\oplus 1\big{]}=\big{\{}(1-\lambda)\ket{+}\!\!\bra{+}\oplus\lambda\colon\lambda\in[0,1]\big{\}}$
    is a generatrix of this cone. One proves along the lines of Ex. 1.1 that the orthogonal projection
    $\pi:\mathcal{D}(\mathrm{M}_{2}\oplus\mathrm{M}_{1})\to\mathcal{D}(\mathcal{R})$
    is not open at any point in the half-open segment
    $\mathcal{G}_{0}:=\big{(}\ket{+}\!\!\bra{+}\oplus 0,0\oplus 1\big{]}=\mathcal{G}\setminus\{\ket{+}\!\!\bra{+}\oplus 0\}\,.$
    Thereby, the equivalence of states and density matrices is described in Coro. 2.7. It is important to observe that $\frac{1}{2}(\mathds{1}_{2}+\cos(\theta)X+\sin(\theta)Z)\oplus 0$ and $\cos(\theta)(X\oplus 1)+\sin(\theta)Z\oplus 0$ are matrices in $\mathrm{M}_{2}\oplus\mathrm{M}_{1}$ for all $\theta\in\mathbb{R}$. The map $\pi$ is open at every point in the complement of $\mathcal{G}_{0}$ because $\mathcal{D}(\mathrm{M}_{2}\oplus\mathrm{M}_{1})$ is a cone over a ball, see [53, Lemma 4.17] for a detailed proof.
  2. b)
    The orthogonal projection $\pi:\mathcal{D}(\mathrm{M}_{2}\oplus\mathrm{M}_{1})\to\mathcal{D}(\mathcal{R})$ has an instructive geometry. The set of density matrices $\mathcal{D}(\mathcal{R})=\pi\big{(}\mathcal{D}(\mathrm{M}_{2})\oplus_{\mathrm{c}}\{1\}\big{)}$ is the convex hull of the projected ball $\pi(\mathcal{D}(\mathrm{M}_{2})\oplus\{0\})$ and the singleton $\pi(0\oplus 1)$ by Lemma 2.5. In turn, $\pi(\mathcal{D}(\mathrm{M}_{2})\oplus\{0\})$ is the filled ellipse of all points
[…]
    $\textstyle=\frac{1}{2}\big{(}\mathds{1}_{3}-M+c_{X}(3M-\mathds{1}_{3})+c_{Z}Z\oplus 0\big{)}\,,$
    where $c_{X},c_{Y},c_{Z}\in\mathbb{R}$ satisfy $c_{X}^{2}+c_{Y}^{2}+c_{Z}^{2}=1$, and $M:=\frac{1}{2}(\ket{+}\!\!\bra{+}\oplus 1)$ is the midpoint of the generatrix $\mathcal{G}$. Note that $\mathds{1}_{3}$, $(3X-\mathds{1}_{2})\oplus 2$, $Z\oplus 0$ is an orthogonal basis of $\mathcal{R}$. The choice of $c_{X}=1,c_{Y}=c_{Z}=0$ yields
    $\pi(\ket{+}\!\!\bra{+}\oplus 0)=M=\pi(0\oplus 1)\,.$ (3.3)
    Thus, $\mathcal{G}$ is perpendicular to $\mathcal{R}$, and $\mathcal{D}(\mathcal{R})=\pi(\mathcal{D}(\mathrm{M}_{2})\oplus\{0\})$ is an ellipse. Moreover, $\pi^{-1}(M)=\mathcal{G}$ holds by equation (3.2) and (3.3).
  3. c)
    The ellipse $\mathcal{D}(\mathcal{R})$ has the semiaxes $\sqrt{3/8}$ and $1/\sqrt{2}$. This follows from the formula (3.2) when $(c_{X},c_{Y},c_{Z})$ is assigned the values of (±$(\pm 1,0,0)$ and $(0,0,\pm 1)$. In retrospect to Rem. 2.9, the value $(1/3,0,\sqrt{8/9})$ confirms that $(\mathds{1}_{3}+\lambda\,Z\oplus 0)/3$ is contained in $\mathcal{D}(\mathcal{R})$ if and only if $|\lambda|\leq\sqrt{2}$.
It is instructive to look at Ex. 3.5 from the perspective of a real *-subalgebra of $\mathrm{M}_{3}$, whose state space can be visualized in three-space.
Example 3.6.
The set of density matrices $\mathcal{D}(\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R}))$ of the real *-subalgebra $\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R})$ of $\mathrm{M}_{3}$ is a symmetric cone, which is the direct convex sum of the great disk $\mathcal{D}(\mathrm{M}_{2}(\mathbb{R}))$ of the Bloch ball and a singleton. As in Ex. 3.5, the closed segment $\mathcal{G}$ is a generatrix of this cone, which is the fiber of the orthogonal projection
$\pi:\mathcal{D}(\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R}))\to\mathcal{D}(\mathcal{R})$
over $M$. The map $\pi$ is not open at any point in $\mathcal{G}_{0}=\mathcal{G}\setminus\{\ket{+}\!\!\bra{+}\oplus 0\}$ and open at every point in the complement of $\mathcal{G}_{0}$. The lack of openness can be visualized graphically in three-space by observing that $\pi$ projects the cone $\mathcal{D}(\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R}))$ along its generatrix $\mathcal{G}$ to the ellipse $\mathcal{D}(\mathcal{R})$.
Lemma 3.7.
Let $\mathcal{R}_{i},\mathcal{R}_{i}^{\prime}$ be real matrix systems on $\mathbb{C}^{n_{i}}$ such that $\mathcal{R}_{i}\subset\mathcal{R}_{i}^{\prime}$, let $\pi_{i}:\mathcal{R}_{i}^{\prime}\to\mathcal{R}_{i}$ denote the orthogonal projection, $i=1,\ldots,m$, and consider the direct sums $\mathcal{R}^{\prime}:=\bigoplus_{i=1}^{m}\mathcal{R}_{i}^{\prime}$ and $\mathcal{R}:=\bigoplus_{i=1}^{m}\mathcal{R}_{i}$. The orthogonal projection $\pi:\mathcal{D}(\mathcal{R}^{\prime})\to\mathcal{D}(\mathcal{R})$ equals
$\pi_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}\pi_{m}:\mathcal{D}(\mathcal{R}_{1}^{\prime})\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}\mathcal{D}(\mathcal{R}_{m}^{\prime})\to\mathcal{D}(\mathcal{R}_{1})\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}\mathcal{D}(\mathcal{R}_{m})\,.$
Proof.
The orthogonal projections $\pi_{i}:\mathcal{D}(\mathcal{R}_{i}^{\prime})\to\mathcal{D}(\mathcal{R}_{i})$, $i=1,\ldots,m$, and $\pi:\mathcal{D}(\mathcal{R}^{\prime})\to\mathcal{D}(\mathcal{R})$ are well defined by Lemma 2.5. A straight-forward computation shows that the orthogonal projection $\pi:\mathcal{R}^{\prime}\to\mathcal{R}$ is the direct sum $\bigoplus_{i=1}^{m}\pi_{i}$. The claim then follows from the third identity of Lemma 3.4 and from the definition of the direct convex sum of maps in formula (3.1). ∎
4. Numerical ranges
The orthogonal projection $\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R})$ may be restricted to the set $\operatorname{ex}\mathcal{D}(\mathrm{M}_{n})$ of extreme points [^1] An point in a convex set $K$ is an extreme point [45] of $K$ if there is no way to express it as a convex combination $(1-s)x+sy$ such that $x,y\in K$ and $s\in(0,1)$, except by taking $x=y$. We denote the set of extreme points of $K$ by $\operatorname{ex}K$. of $\mathcal{D}(\mathrm{M}_{n})$. The openness of this restriction was studied in matrix theory [17, 29, 30, 31, 34, 35, 49, 50, 54] for two-dimensional state spaces $\mathcal{D}(\mathcal{R})$ represented as numerical ranges (see Rem. 4.3).
Let $A_{1},\ldots,A_{k}\in\mathcal{H}(\mathrm{M}_{n})$ and consider the real matrix system
$\mathcal{R}(A_{1},\ldots,A_{k}):=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{n},A_{1},\ldots,A_{k})\,.$
The image of $\mathcal{D}(\mathrm{M}_{n})$ under the real linear map
$v:\mathcal{H}(\mathrm{M}_{n})\to\mathbb{R}^{k}\,,\quad B\mapsto(\langle B,A_{1}\rangle,\dots,\langle B,A_{k}\rangle){}^{\mathrm{T}}$
is the joint numerical range $V(A_{1},\ldots,A_{k}):=v(\mathcal{D}(\mathrm{M}_{n}))\subset\mathbb{R}^{k}$, see [9]. Aware of the notational imprecision, we use the same label $v$ also for several restrictions of $v$, among others for
$v:\mathcal{D}(\mathrm{M}_{n})$ $\to V(A_{1},\ldots,A_{k})\,,$ $\rho$ $\mapsto(\langle\rho,A_{1}\rangle,\dots,\langle\rho,A_{k}\rangle){}^{\mathrm{T}}\,.$ (4.1)
Lemma 4.1.
The following diagram commutes.
${\mathcal{D}(\mathrm{M}_{n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$$\scriptstyle{v}$${V(A_{1},\ldots,A_{k})}$${\mathcal{D}(\mathcal{R}(A_{1},\ldots,A_{k}))}$$\scriptstyle{v}$
Proof.
Let $\mathcal{R}:=\mathcal{R}(A_{1},\ldots,A_{k})$. By Lemma 2.5, the orthogonal projection $\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R})$ is surjective. It is straightforward to verify that $v:\mathcal{H}(\mathrm{M}_{n})\to\mathbb{R}^{k}$ factors through $\mathcal{R}$, in the sense that $v=v\circ\pi$ holds. Hence, the commutativity of the diagram is implied by the injectivity of $v$ restricted to the affine space $\{B\in\mathcal{H}(\mathcal{R}):\operatorname{tr}(B)=1\}$. Let $B_{1},B_{2}$ be contained in this affine space. If $v(B_{1})=v(B_{2})$, then
00 0 $=v(B_{1}-B_{2})=\langle B_{1}-B_{2},A_{i}\rangle_{i=1}^{k}$
[…]
This implies $B_{1}-B_{2}=0$ as $B_{1}-B_{2}\in\mathcal{H}(\mathcal{R})=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{n},A_{1},\ldots,A_{k})$. ∎
In the remainder of this section, let $k=2$ and $A:=A_{1}+\mathrm{i}A_{2}$. The image of the unit sphere $\mathbb{C}S^{n}:=\{\ket{\varphi}\in\mathbb{C}^{n}:\braket{\varphi}{\varphi}=1\}$ under the hermitian quadratic form $f_{A}:\mathbb{C}^{n}\to\mathbb{C}$, |$\ket{\varphi}\mapsto\braket{\varphi}{A\varphi}$ is the numerical range
$W(A):=f_{A}(\mathbb{C}S^{n})\subset\mathbb{C}\,.$
Here, $\braket{\varphi_{1}}{\varphi_{2}}:=\overline{x_{1}}y_{1}+\cdots+\overline{x_{n}}y_{n}$ is the inner product of |$\ket{\varphi_{1}}=(x_{1},\ldots,x_{n}){}^{\mathrm{T}}$ and |$\ket{\varphi_{2}}=(y_{1},\ldots,y_{n}){}^{\mathrm{T}}$ in $\mathbb{C}^{n}$. We use the same label $f_{A}$ to denote the map
$f_{A}:\mathbb{C}S^{n}\to W(A)\,,\quad\ket{\varphi}\mapsto\braket{\varphi}{A\varphi}\,.$
Minkowski’s theorem [47] asserts that every compact convex set is the convex hull of its extreme points.
Proposition 4.2.
The following diagram commutes.
${\mathbb{C}S^{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$|$\scriptstyle{\ket{\varphi}\mapsto\ket{\varphi}\!\bra{\varphi}}$ |$\scriptstyle{f_{A}}$${\operatorname{ex}\mathcal{D}(\mathrm{M}_{n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{v}$$\scriptstyle{\rho\mapsto\rho}$${\mathcal{D}(\mathrm{M}_{n})}$$\scriptstyle{v}$$\scriptstyle{\pi}$${W(A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${V(A_{1},A_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$z\mapsto\left(\begin{smallmatrix}\mathrm{Re}(z)\\ \mathrm{Im}(z)\end{smallmatrix}\right)$${\mathcal{D}(\mathcal{R}(A_{1},A_{2}))}$$\scriptstyle{v}$
Proof.
The bottom right triangle is the case $k=2$ of Lemma 4.1. The map $f_{A}:\mathbb{C}S^{n}\to W(A)$ factors through $\operatorname{ex}\mathcal{D}(\mathrm{M}_{n})$, $\mathcal{D}(\mathrm{M}_{n})$, and $V(A_{1},A_{2})$, as $\mathbb{C}S^{n}\to\mathcal{D}(\mathrm{M}_{n})$, |$\ket{\varphi}\mapsto\ket{\varphi}\!\!\bra{\varphi}$ | parametrizes the extreme points of $\mathcal{D}(\mathrm{M}_{n})$, see for example [7, Sec. 5.1], and since for all |$\ket{\varphi}\in\mathbb{C}S^{n}$ we have
[…]
It remains to show that $g:W(A)\to V(A_{1},A_{2})$, $z\mapsto(\mathrm{Re}(z),\mathrm{Im}(z)){}^{\mathrm{T}}$ is onto (being the restriction of a bijection, the map $g$ is one-to-one).
First, we show that $\operatorname{ex}V(A_{1},A_{2})$ is included in the image of $g$. The preimage of every extreme point $x$ of $V(A_{1},A_{2})$ under $v:\mathcal{D}(\mathrm{M}_{n})\to V(A_{1},A_{2})$ contains an extreme point of $\mathcal{D}(\mathrm{M}_{n})$. This is true because the preimage of $x$ is a face $\mathcal{F}$ of $\mathcal{D}(\mathrm{M}_{n})$, which has an extreme point $\rho$ by Minkowski’s theorem, since $\mathcal{D}(\mathrm{M}_{n})$, and hence $\mathcal{F}$, is compact. Since $\rho$ is also an extreme point of $\mathcal{D}(\mathrm{M}_{n})$, the claim follows from $\mathbb{C}S^{n}\to\operatorname{ex}\mathcal{D}(\mathrm{M}_{n})$ being onto.
Second, the convex hull of $\operatorname{ex}V(A_{1},A_{2})$ is included in the image of $g$ because $W(A)$ is convex by the Toeplitz-Hausdorff theorem [19, 37]. Third, the map $g$ is onto, because $V(A_{1},A_{2})$ is the convex hull of its extreme points, again by Minkowski’s theorem. ∎
The affine isomorphism $W(A)\cong\mathcal{D}(\mathcal{R}(A_{1},A_{2}))$ of Prop. 4.2 has an analogue in the much more general setting of matrix-valued states [22, Thm. 5.1].
Let $f_{A}^{-1}$ denote the multi-valued inverse of $f_{A}:\mathbb{C}S^{n}\to W(A)$. Corey et al. [17] define $f_{A}^{-1}$ to be strongly continuous at $z\in W(A)$ if the map $f_{A}$ is open at every point in the fiber $f_{A}^{-1}(z)$ of $f_{A}$ over $z$.
Remark 4.3.
Strong continuity can be described in terms of standard results on the numerical range. We refer to Sec. 8 of [29] and the references therein.
There exists a family of orthonormal bases |$\ket{\varphi_{1}(\theta)},\dots,\ket{\varphi_{n}(\theta)}$ of $\mathbb{C}^{n}$ that is analytically parametrized by a real number $\theta\in\mathbb{R}$; and there are analytic functions $\lambda_{i}:\mathbb{R}\to\mathbb{R}$, called eigenfunctions [30], such that
$(\cos(\theta)A_{1}+\sin(\theta)A_{2})\ket{\varphi_{i}(\theta)}=\lambda_{i}(\theta)\ket{\varphi_{i}(\theta)}\,,\quad\theta\in\mathbb{R}\,,\quad i=1,\dots,n\,.$
For every $i=1,\ldots,n$, the numerical range $W(A)$ includes the image $\operatorname{Img}(z_{i})$ of the curve
$z_{i}:\mathbb{R}\to\mathbb{C}\,,\quad\theta\mapsto e^{\mathrm{i}\theta}(\lambda_{i}(\theta)+\mathrm{i}\lambda^{\prime}_{i}(\theta))\,.$
Every extreme point of $W(A)$ is contained in $\operatorname{Img}(z_{i})$ for some $i=1,\ldots,n$.
If $z\in W(A)$ is not an extreme point of $W(A)$, then $f_{A}^{-1}$ is strongly continuous at $z$ [17, Thm. 4]. Let $z$ be an extreme point of $W(A)$. Then there are $\theta_{0}\in\mathbb{R}$ and $i_{0}\in\{1,\ldots,n\}$ such that $z=z_{i_{0}}(\theta_{0})$. Now, $f_{A}^{-1}$ is strongly continuous at $z$ if and only if $z_{i}(\theta_{0})=z$ implies $z_{i}=z_{i_{0}}$ for all $i=1,\ldots,n$, see [30, Thm. 2.1.1]. Leake at al. [30] state the latter condition by saying that the eigenfunctions corresponding to $z$ at $\theta_{0}$ do not split.
Remark 4.4.
Strong continuity is connected to the openness of a linear map. For all $z\in W(A)$, the multi-valued map $f_{A}^{-1}$ is strongly continuous at $z$ if and only if the restricted linear map $v:\mathcal{D}(\mathrm{M}_{n})\to V(A_{1},A_{2})$, introduced in (4.1), is open at every point in the fiber $v^{-1}[(\mathrm{Re}(z),\mathrm{Im}(z)){}^{\mathrm{T}}]$, see [54, Coro. 5.2]. Moreover, for every $x\in V(A_{1},A_{2})$, the map $v$ is open at some point in the relative interior [^2] The relative interior [45] of a convex set $K$, denoted by $\operatorname{ri}(K)$, is the interior of $K$ in the relative topology of the affine hull of $K$. of $v^{-1}(x)$ if and only if $v$ is open at every point in $v^{-1}(x)$.
Proposition 4.5.
Let $z\in W(A)$, let $x:=(\mathrm{Re}(z),\mathrm{Im}(z)){}^{\mathrm{T}}\in\mathbb{R}^{2}$, let $\rho$ be the unique density matrix of $\mathcal{R}(A_{1},A_{2})$ that satisfies $v(\rho)=x$, and let $\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R}(A_{1},A_{2}))$ denote the orthogonal projection. Then the following assertions are equivalent.
  • $f_{A}^{-1}$ is strongly continuous at $z$,
  • $v:\mathcal{D}(\mathrm{M}_{n})\to V(A_{1},A_{2})$ is open at every point in $v^{-1}(x)$,
  • $\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R}(A_{1},A_{2}))$ is open at every point in $\pi^{-1}(\rho)$.
We have $z\in\operatorname{ex}W(A)\Leftrightarrow x\in\operatorname{ex}V(A_{1},A_{2})\Leftrightarrow\rho\in\operatorname{ex}\mathcal{D}(\mathcal{R}(A_{1},A_{2}))$. If $z\in\operatorname{ex}W(A)$, then $z=z_{i_{0}}(\theta_{0})$ for some $\theta_{0}\in\mathbb{R}$ and $i_{0}\in\{1,\ldots,n\}$, and the following assertions are equivalent.
  • $f_{A}^{-1}$ is strongly continuous at $z$,
  • the eigenfunctions corresponding to $z$ at $\theta_{0}$ do not split,
  • $\pi$ is open at some point in the relative interior of $\pi^{-1}(\rho)$.
Proof.
The claims follow directly from Rem. 4.3 and 4.4, and Prop. 4.2. ∎
Example 4.6.
Prop. 4.5 ignores that $\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R}(A_{1},A_{2}))$ could be open at some point in the fiber $\pi^{-1}(\rho)$ over an extreme point $\rho$ of $\mathcal{D}(\mathcal{R}(A_{1},A_{2}))$, but not open anywhere in the relative interior of $\pi^{-1}(\rho)$.
This occurs for $A_{1}:=X\oplus 1$ and $A_{2}:=Z\oplus 0$. As discussed in Ex. 3.5 a) and b), the orthogonal projection $\pi_{\mathcal{R}}:\mathcal{D}(\mathrm{M}_{2}\oplus\mathrm{M}_{1})\to\mathcal{D}(\mathcal{R}(A_{1},A_{2}))$ is open at |++|$\ket{+}\!\!\bra{+}\oplus 0$ and nowhere else in the fiber $\pi_{\mathcal{R}}^{-1}(M)=\big{[}\ket{+}\!\!\bra{+}\oplus 0,0\oplus 1\big{]}$ over $M=\frac{1}{2}(\ket{+}\!\!\bra{+}\oplus 1)$. In particular, $\pi_{\mathcal{R}}$ is not open anywhere in the relative interior of $\pi_{\mathcal{R}}^{-1}(M)$. Ex. 8.3 proves the analogue for the larger fiber $\pi^{-1}(M)$.
5. Retractions
Generalizing a known result, we prove that the set of density matrices of every real *-subalgebra of $\mathrm{M}_{n}$ is a stable convex set. The proof relies on retractions of state spaces, a topic that also proved helpful in the foundations of quantum information theory [24].
Remark 5.1 (Stability of state spaces).
  1. a)
    The stability problem of a finite-dimensional compact convex set $K$ is completely solved [40]. The $d$-skeleton of $K$ is defined as the union of all faces [^3] A face [45] of a convex set $K$ is a convex subset $F\subset K$ such that $x,y\in F$ is implied by $(1-s)x+sy\in F$ for all $x,y\in K$ and $s\in(0,1)$. of $K$ of dimension at most $d$. The convex set $K$ is stable if and only if for every nonnegative integer $d$, the $d$-skeleton of $K$ is closed.
  2. b)
    The set of density matrices $\mathcal{D}(\mathrm{M}_{n})$ is stable [20] because all its $d$-skeletons are closed. The closedness follows from three arguments: First, the set $\mathcal{D}(\mathrm{M}_{n})$ is a compact convex set of dimension $n^{2}-1$. Second, every nonempty face of $\mathrm{M}_{n}$ is unitarily similar to $\mathcal{D}(\mathrm{M}_{l})\oplus\{0\}$ for some positive integer $l$, and third, the unitary group $U(n)$ is compact.
  3. c)
    The state space $\mathcal{S}(\mathcal{A})$ of every *-subalgebra $\mathcal{A}$ of $\mathrm{M}_{n}$ is stable [41] as it is a direct convex sum of state spaces of full matrix algebras $\mathrm{M}_{n}$. As $\mathcal{S}(\mathcal{A})$ is stable, the set of density matrices $\mathcal{D}(\mathcal{A})$ is stable, too, by Lemma 2.2.
A retraction is an affine map $f:K\to L$ between compact convex sets $K,L$ which is left-inverse to an affine map $g:L\to K$, called a section.
Proposition 5.2.
If $\mathcal{A}$ is a real *-subalgebra of $\mathrm{M}_{n}$, then the orthogonal projection $\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{A})$ is a retraction and $\mathcal{D}(\mathcal{A})$ is stable.
Proof.
The orthogonal projection $\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{A})$ is well defined by Lemma 2.5 and it is a retraction because the inclusion $\mathcal{D}(\mathcal{A})\subset\mathcal{D}(\mathrm{M}_{n})$ of equation (2.4) provides a section for $\pi$. Coro. 1.3 in [15] asserts that the image of a stable convex set under a retraction is stable. Therefore, and since $\mathcal{D}(\mathrm{M}_{n})$ is stable by Rem. 5.1 b), the convex set $\mathcal{D}(\mathcal{A})$ is stable. ∎
Example 5.3.
Prop. 5.2 does not generalize to arbitrary matrix systems. We consider the hermitian matrices
$A_{1}:=X\oplus 1\oplus 1\,,\quad A_{2}:=Z\oplus 0\oplus 0\,,\quad A_{3}:=0\oplus(-1)\oplus 1\,,$
in the algebra $\mathrm{M}_{2}\oplus\mathrm{M}_{1}\oplus\mathrm{M}_{1}$. The joint numerical range $V(A_{1},A_{2},A_{3})$, introduced in Sec. 4, is easier to handle algebraically than the set of density matrices $\mathcal{D}(\mathcal{R}(A_{1},A_{2},A_{3}))$, to which it is affinely isomorphic by Lemma 4.1. Since $A_{1},A_{2},A_{3}\in\mathrm{M}_{2}\oplus\mathrm{M}_{1}\oplus\mathrm{M}_{1}$, we have the standard result of
$V(A_{1},A_{2},A_{3})=\operatorname{conv}\big{(}V(X,Z,0)\cup V(1,0,-1)\cup V(1,0,1)\big{)}\,.$
Hence, $V(A_{1},A_{2},A_{3})=\operatorname{conv}(S)$ is the convex hull of
$S:=\{(c_{1},c_{2},0){}^{\mathrm{T}}\in\mathbb{R}^{3}\mid c_{1}^{2}+c_{2}^{2}=1\}\cup\{(1,0,-1){}^{\mathrm{T}},(1,0,1){}^{\mathrm{T}}\}\,.$
The set of extreme points $S\setminus\{(1,0,0){}^{\mathrm{T}}\}$ of $V(A_{1},A_{2},A_{3})$ is not closed, hence $V(A_{1},A_{2},A_{3})$ is not stable by Rem. 5.1 a).
6. Proof of Thm. 1.2
The main ideas in establishing Thm. 1.2 are that the set of density matrices $\mathcal{D}(\mathrm{M}_{n})$ is a stable convex set and that $\mathcal{D}(\mathrm{M}_{n})$ is highly symmetric. Being stable and symmetric, $\mathcal{D}(\mathrm{M}_{p+q})$ projects openly onto $\mathcal{D}(\mathrm{M}_{p}\oplus\mathrm{M}_{q})$. Loosely speaking, Thm. 1.2 is obtained by combining various such open projections, since every *-subalgebra of $\mathrm{M}_{n}$ is a direct sum of full matrix algebras.
If $K$ is a stable convex set, then the arithmetic mean map
$K^{\times k}\to K\,,\quad(x_{1},\dots,x_{k})\mapsto\frac{1}{k}\sum_{i=1}^{k}x_{i}\,,$
defined on the $k$-fold cartesian product of $K$, is open for all positive integers $k$. More generally, for every tuple $(s_{1},\dots,s_{k})$ of nonnegative real numbers adding up to one, the map
$K^{\times k}\to K\,,\quad(x_{1},\dots,x_{k})\mapsto\sum_{i=1}^{k}s_{i}x_{i}$
is open. The latter assertion is proved for $k=2$ in Prop. 1.1 in [16]. By induction, it is true for every $k>2$ as well.
Lemma 6.1.
Let $\mathcal{R}$ be a real matrix system on $\mathbb{C}^{n}$ and let $\mathcal{D}(\mathcal{R})$ be stable and invariant under an orthogonal transformation $\gamma:\mathcal{R}\to\mathcal{R}$ that generates a finite cyclic group. Let $\mathcal{A}:=\{A\in\mathcal{R}:\gamma(A)=A\}$ be a real *-subalgebra of $\mathrm{M}_{n}$. Then the orthogonal projection $\mathcal{D}(\mathcal{R})\to\mathcal{D}(\mathcal{A})$ is open.
Proof.
Let $k$ denote the order of the cyclic group generated by $\gamma$. The main idea is to write the orthogonal projection $\pi:\mathcal{D}(\mathcal{R})\to\mathcal{D}(\mathcal{A})$ in terms of the arithmetic mean map
$a:\mathcal{D}(\mathcal{R})^{\times k}\to\mathcal{D}(\mathcal{R})\,,\quad(\rho_{1},\dots,\rho_{k})\mapsto\frac{1}{k}\sum_{i=1}^{k}\rho_{i}\,,$
which, as discussed above, is open because $\mathcal{D}(\mathcal{R})$ is stable. The proof is done in two steps. First, we prove that
$a(\gamma(\mathcal{O})\times\dots\times\gamma^{k}(\mathcal{O}))\cap\mathcal{A}$ (6.1)
is an open subset of $\mathcal{D}(\mathcal{A})$ for all open subsets $\mathcal{O}$ of $\mathcal{D}(\mathcal{R})$. Secondly, we prove that
$a(\gamma(\mathcal{K})\times\dots\times\gamma^{k}(\mathcal{K}))\cap\mathcal{A}=\pi(\mathcal{K})$ (6.2)
holds for all convex subsets $\mathcal{K}$ of $\mathcal{D}(\mathcal{R})$. The assertions (6.1) and (6.2) together show that $\pi(\mathcal{O})$ is an open subset of $\mathcal{D}(\mathcal{A})$ for all convex open subsets $\mathcal{O}$ of $\mathcal{D}(\mathcal{R})$, and hence for all open subsets.
First, we prove that the set in (6.1) is open. As $\mathcal{D}(\mathcal{R})$ is invariant under the orthogonal transformation $\gamma$, the map $\gamma$ restricts to a homeomorphism $\mathcal{D}(\mathcal{R})\to\mathcal{D}(\mathcal{R})$. It follows that $\gamma(\mathcal{O})\times\dots\times\gamma^{k}(\mathcal{O})$ is an open subset of the $k$-fold cartesian product $\mathcal{D}(\mathcal{R})^{\times k}$. Then
$\widetilde{\mathcal{O}}:=a(\gamma(\mathcal{O})\times\dots\times\gamma^{k}(\mathcal{O}))$
is an open subset of $\mathcal{D}(\mathcal{R})$, because $\mathcal{D}(\mathcal{R})$ is stable. Finally, $\widetilde{\mathcal{O}}\cap\mathcal{A}$ is an open subset of $\mathcal{D}(\mathcal{A})$ by equation (2.4), which shows that $\widetilde{\mathcal{O}}\cap\mathcal{A}$ equals $\widetilde{\mathcal{O}}\cap\mathcal{D}(\mathcal{A})$.
Secondly, we prove the formula (6.2), beginning with the inclusion $\supset$”. Let $\rho\in\mathcal{K}\subset\mathcal{D}(\mathcal{R})$. The density matrix $\sigma:=a(\gamma(\rho),\dots,\gamma^{k}(\rho))$ is invariant under $\gamma$. This implies $\sigma\in\mathcal{A}$, hence $\sigma\in\mathcal{D}(\mathcal{A})$ by (2.4). Since $\gamma$ is self-adjoint, for all $A\in\mathcal{H}(\mathcal{A})$ we have
$\mathrm{Re}\langle\rho,A\rangle=\mathrm{Re}\langle a(\rho,\dots,\rho),A\rangle=\mathrm{Re}\langle a(\gamma(\rho),\dots,\gamma^{k}(\rho)),A\rangle=\mathrm{Re}\langle\sigma,A\rangle\,,$
hence $\pi(\rho)=\sigma$. Regarding the inclusion $\subset$”, let $\rho_{i}\in\mathcal{K}$, $i=1,\dots,k$. Let $\sigma:=a(\gamma(\rho_{1}),\dots,\gamma^{k}(\rho_{k}))$ and $\tau:=a(\rho_{1},\dots,\rho_{k})$. Then for all $A\in\mathcal{H}(\mathcal{A})$
$\mathrm{Re}\langle\sigma,A\rangle=\mathrm{Re}\langle a(\gamma(\rho_{1}),\dots,\gamma^{k}(\rho_{k})),A\rangle=\mathrm{Re}\langle a(\rho_{1},\dots,\rho_{k}),A\rangle=\mathrm{Re}\langle\tau,A\rangle$
holds. If $\sigma\in\mathcal{A}$, then $\sigma=\pi(\tau)$ follows. As $\mathcal{K}$ is convex, $\tau\in\mathcal{K}$ holds and we obtain $\sigma\in\pi(\mathcal{K})$. ∎
Remark 6.2.
A hermitian matrix $M\in\mathcal{H}(M_{p+q})$ in the block form
$M=\begin{pmatrix}A&B\\ B^{\ast}&C\end{pmatrix}\,,\quad A\in\mathcal{H}(M_{p})\,,\quad B\in\mathbb{C}^{p\times q}\,,\quad C\in\mathcal{H}(M_{q})$
is positive semidefinite if and only if the top left block $A$ is positive semidefinite, the range of $B$ is included in the range of $A$, and the generalized Schur complement $M/A=C-B^{\ast}A^{-}B$ is positive semidefinite, where $A^{-}$ is a generalized inverse of $A$, that is to say, $A^{-}\in M_{p}$ and $AA^{-}A=A$ holds [25].
Proposition 6.3.
The orthogonal projection $\mathcal{D}(\mathrm{M}_{p+q})\to\mathcal{D}(\mathrm{M}_{p}\oplus\mathrm{M}_{q})$ is open.
Proof.
Lemma 6.1 proves the claim when $\mathcal{R}:=\mathrm{M}_{p+q}$ and $\mathcal{A}:=\mathrm{M}_{p}\oplus\mathrm{M}_{q}$. The set of density matrices $\mathcal{D}(\mathrm{M}_{p+q})$ is stable by Rem. 5.1 b). The reflection $\gamma:\mathrm{M}_{p+q}\to\mathrm{M}_{p+q}$ at the subspace $\mathrm{M}_{p}\oplus\mathrm{M}_{q}$ generates a group of order two.
We show that $\mathcal{D}(\mathrm{M}_{p+q})$ is invariant under $\gamma$. In block form, the reflection reads
$\gamma:\begin{pmatrix}A&B\\ D&C\end{pmatrix}\mapsto\begin{pmatrix}A&-B\\ -D&C\end{pmatrix}\,,$
where $A\in\mathrm{M}_{p}$, $B\in\mathbb{C}^{p\times q}$, $C\in\mathrm{M}_{q}$, and $D\in\mathbb{C}^{q\times p}$. The space of hermitian matrices is invariant under $\gamma$, which restricts to
$\mathcal{H}(\mathrm{M}_{p+q})\to\mathcal{H}(\mathrm{M}_{p+q})\,,\quad M=\begin{pmatrix}A&B\\ B^{\ast}&C\end{pmatrix}\mapsto M^{\prime}=\begin{pmatrix}A&-B\\ -B^{\ast}&C\end{pmatrix}\,,$
where $A\in\mathcal{H}(\mathrm{M}_{p})$, $B\in\mathbb{C}^{p\times q}$, and $C\in\mathcal{H}(\mathrm{M}_{q})$. By Rem. 6.2, the map $M\mapsto M^{\prime}$ preserves the positive semidefiniteness since $M$ and $M^{\prime}$ have the same diagonal blocks and both off-diagonal blocks differ in a sign, so that $M^{\prime}/A=M/A$ holds. The map $\gamma$ preserves the trace. The set of fixed points $\mathrm{M}_{p}\oplus\mathrm{M}_{q}=\{A\in\mathrm{M}_{p+q}:\gamma(A)=A\}$ is a *-subalgebra of $\mathrm{M}_{p+q}$. ∎
Corollary 6.4.
The orthogonal projection $\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathrm{M}_{n_{1}}\oplus\dots\oplus\mathrm{M}_{n_{k}})$ is open, where $n=n_{1}+\dots+n_{k}$.
Proof.
Proceeding by induction, we observe that the orthogonal projection $\mathcal{D}(\mathrm{M}_{n_{1}})\to\mathcal{D}(\mathrm{M}_{n_{1}})$ is of course open. Let $k\geq 1$ and assume that the orthogonal projection
$\pi_{k}:\mathcal{D}(\mathrm{M}_{n_{1}+\dots+n_{k}})\to\mathcal{D}(\mathrm{M}_{n_{1}}\oplus\dots\oplus\mathrm{M}_{n_{k}})$
is open. The orthogonal projection $\mathrm{M}_{n_{1}+\dots+n_{k+1}}\to\mathrm{M}_{n_{1}}\oplus\dots\oplus\mathrm{M}_{n_{k+1}}$ factors into the orthogonal projections
$\mathrm{M}_{n_{1}+\dots+n_{k+1}}$ $\to\mathrm{M}_{n_{1}+\dots+n_{k}}\oplus\mathrm{M}_{n_{k+1}}$
$\text{and}\quad\mathrm{M}_{n_{1}+\dots+n_{k}}\oplus\mathrm{M}_{n_{k+1}}$ $\to\mathrm{M}_{n_{1}}\oplus\dots\oplus\mathrm{M}_{n_{k}}\oplus\mathrm{M}_{n_{k+1}}\,.$
Hence, by Lemma 2.5, the map $\pi_{k+1}$ factors into the orthogonal projections
$\mathcal{D}(\mathrm{M}_{n_{1}+\dots+n_{k+1}})$ $\to\mathcal{D}(\mathrm{M}_{n_{1}+\dots+n_{k}}\oplus\mathrm{M}_{n_{k+1}})$ (6.3)
$\text{and}\quad\mathcal{D}(\mathrm{M}_{n_{1}+\dots+n_{k}}\oplus\mathrm{M}_{n_{k+1}})$ $\to\mathcal{D}(\mathrm{M}_{n_{1}}\oplus\dots\oplus\mathrm{M}_{n_{k}}\oplus\mathrm{M}_{n_{k+1}})\,.$ (6.4)
The map (6.3) is open by Prop. 6.3. Lemma 3.7 shows that (6.4) equals
$\mathcal{D}(\mathrm{M}_{n_{1}+\dots+n_{k}})\oplus_{\mathrm{c}}\mathcal{D}(\mathrm{M}_{n_{k+1}})$ $\to\mathcal{D}(\mathrm{M}_{n_{1}}\oplus\dots\oplus\mathrm{M}_{n_{k}})\oplus_{\mathrm{c}}\mathcal{D}(\mathrm{M}_{n_{k+1}})\,,$
which is open by Prop. 3.2 (for $m=2$) and by the induction hypothesis. Being the composition of two open maps, $\pi_{k+1}$ is open. ∎
Proposition 6.5.
The orthogonal projection $\mathcal{D}(\bigoplus_{i=1}^{k}\mathrm{M}_{q})\to\mathcal{D}(\mathrm{M}_{q}\otimes\mathds{1}_{k})$ is open.
Proof.
Lemma 6.1 proves the claim when $\mathcal{R}:=\bigoplus_{i=1}^{k}\mathrm{M}_{q}$ and $\mathcal{A}:=\mathrm{M}_{q}\otimes\mathds{1}_{k}$. The convex set $\mathcal{D}(\mathcal{R})$ is stable by Rem. 5.1 c) and equals the $k$-fold direct convex sum $\mathcal{D}(\mathcal{R})=\mathcal{D}(\mathrm{M}_{q})\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}\mathcal{D}(\mathrm{M}_{q})$ of $\mathcal{D}(\mathrm{M}_{q})$ by Lemma 3.4. Hence, the cyclic permutation $\gamma=(1,\dots,k)$ defines the orthogonal transformation
$\gamma:\mathcal{D}(\mathcal{R})\to\mathcal{D}(\mathcal{R})\,,\quad(\sigma_{1},\dots,\sigma_{k})\mapsto(\sigma_{\gamma^{-1}(1)},\dots,\sigma_{\gamma^{-1}(k)})\,,$
which generates a group of order $k$. Clearly, $\mathcal{D}(\mathcal{R})$ is invariant under $\gamma$ and $\mathcal{A}=\{A\in\mathcal{R}:\gamma(A)=A\}$ is a *-subalgebra of $\mathrm{M}_{kq}$. ∎
Proof of Thm. 1.2.
Since $\mathcal{A}$ is a *-subalgebra of $\mathrm{M}_{n}$, there exists a unitary $n\times n$ matrix $U$ such that $U\mathcal{A}U^{\ast}=\bigoplus_{i=1}^{m}\mathcal{A}_{i}$, where $\mathcal{A}_{i}:=\mathrm{M}_{q_{i}}\otimes\mathds{1}_{k_{i}}$ for every $i=1,\dots,m$, and $q_{1}k_{1}+\dots+q_{m}k_{m}=n$, see [21, Thm. 5.6].
As $\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathrm{M}_{n})$, $\rho\mapsto U\rho U^{\ast}$ is a homeomorphism, it suffices to prove that the orthogonal projection $\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\bigoplus_{i=1}^{m}\mathcal{A}_{i})$ is open. The orthogonal projection $\mathrm{M}_{n}\to\bigoplus_{i=1}^{m}\mathcal{A}_{i}$ factors into the orthogonal projections
$\mathrm{M}_{n}$ $\textstyle\to\mathcal{B}_{1}\oplus\dots\oplus\mathcal{B}_{m}$
$\text{and}\quad\mathcal{B}_{1}\oplus\dots\oplus\mathcal{B}_{m}$ $\textstyle\to\mathcal{A}_{1}\oplus\dots\oplus\mathcal{A}_{m}\,,$
where $\mathcal{B}_{i}:=\bigoplus_{j=1}^{k_{i}}\mathrm{M}_{q_{i}}$, $i=1,\dots,m$. By Lemma 2.5, the map $\pi$ factors into
$\mathcal{D}(\mathrm{M}_{n})$ $\to\mathcal{D}(\mathcal{B}_{1}\oplus\dots\oplus\mathcal{B}_{m})\,,$ (6.5)
$\text{and}\quad\mathcal{D}(\mathcal{B}_{1}\oplus\dots\oplus\mathcal{B}_{m})$ $\to\mathcal{D}(\mathcal{A}_{1}\oplus\dots\oplus\mathcal{A}_{m})\,.$ (6.6)
The map (6.5) is open by Coro. 6.4 (for $k=k_{1}+\dots+k_{m}$). Lemma 3.7 shows that the map (6.6) equals
$\mathcal{D}(\mathcal{B}_{1})\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}\mathcal{D}(\mathcal{B}_{m})$ $\to\mathcal{D}(\mathcal{A}_{1})\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}\mathcal{D}(\mathcal{A}_{m})\,,$
which is open by Prop. 3.2 and Prop. 6.5. In conclusion, $\pi$ is open as it is a composition of two open maps. ∎
7. Real *-subalgebras of $\mathrm{M}_{n}$
Every real *-subalgebra of $\mathrm{M}_{n}$ is *-isomorphic [21, Thm. 5.22] to a direct sum of algebras of real, complex, and quaternionic $q$-by-$q$-matrices of various sizes $q$. We are here interested in the algebra $\mathrm{M}_{n}(\mathbb{R})$ of real $n\times n$ matrices.
Proposition 7.1.
The orthogonal projection $\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathrm{M}_{n}(\mathbb{R}))$ is open.
Proof.
Lemma 6.1 proves the claim when $\mathcal{R}:=\mathrm{M}_{n}$ and $\mathcal{A}:=\mathrm{M}_{n}(\mathbb{R})$. The set of density matrices $\mathcal{D}(\mathrm{M}_{n})$ is stable by Rem. 5.1 b). The reflection $\gamma:\mathrm{M}_{n}\to\mathrm{M}_{n}$ at the real subspace $\mathrm{M}_{n}(\mathbb{R})$ generates a group of order two.
The algebra $\mathrm{M}_{n}$ is the orthogonal direct sum $\mathrm{M}_{n}=\mathrm{M}_{n}(\mathbb{R})\oplus\mathrm{i}\,\mathrm{M}_{n}(\mathbb{R})$ and the reflection at the real subspace $\mathrm{M}_{n}(\mathbb{R})$ reads
$\gamma:\mathrm{M}_{n}\to\mathrm{M}_{n}\,,\quad A+\mathrm{i}B\mapsto A-\mathrm{i}B\,,\quad A,B\in\mathrm{M}_{n}(\mathbb{R})\,.$
The orthogonal transformation $\gamma$ preserves the space of hermitian matrices, which is the orthogonal direct sum
$\mathcal{H}(\mathrm{M}_{n})=\mathrm{Sym}_{n}(\mathbb{R})\oplus\mathrm{i}\,\mathrm{Skew}_{n}(\mathbb{R})$
of the space of real symmetric matrices
$\mathrm{Sym}_{n}(\mathbb{R})$ :=$:=\{A\in\mathrm{M}_{n}(\mathbb{R})\colon A{}^{\mathrm{T}}=A\}=\mathcal{H}(\mathrm{M}_{n}(\mathbb{R}))$
and the space of skew-symmetric matrices
$\mathrm{Skew}_{n}(\mathbb{R})$ :=$:=\{A\in\mathrm{M}_{n}(\mathbb{R})\colon A{}^{\mathrm{T}}=-A\}\,.$
We prove that $\gamma$ preserves the trace and the positive semidefiniteness on the space of hermitian matrices. Let $A,C\in\mathrm{Sym}_{n}(\mathbb{R})$ and $B,D\in\mathrm{Skew}_{n}(\mathbb{R})$. Then $\operatorname{tr}(A+\mathrm{i}B)=\operatorname{tr}(A)$ shows that the trace is preserved. It is well known that a matrix is positive semidefinite if and only if its inner product with the square of every hermitian matrix is nonnegative. Thus
$\langle A-\mathrm{i}B,(C+\mathrm{i}D)^{2}\rangle$ $=\langle A+\mathrm{i}B,(C-\mathrm{i}D)^{2}\rangle$
shows that $A-\mathrm{i}B$ is positive semidefinite if $A+\mathrm{i}B$ is positive semidefinite. Clearly, $\mathrm{M}_{n}(\mathbb{R})=\{A\in\mathrm{M}_{n}:\gamma(A)=A\}$ is a real *-subalgebra of $\mathrm{M}_{n}$. ∎
As the orthogonal projection $\mathrm{M}_{n}\to\mathrm{M}_{n}(\mathbb{R})$ is the entrywise real part, we denote by $\mathrm{Re}:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathrm{M}_{n}(\mathbb{R}))$ the orthogonal projection of $\mathcal{D}(\mathrm{M}_{n})$ onto $\mathcal{D}(\mathrm{M}_{n}(\mathbb{R}))$.
Example 7.2.
We consider the chain $\mathrm{M}_{3}\supset\mathrm{M}_{2}\oplus\mathrm{M}_{1}\supset\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R})$ of real *-subalgebras of $\mathrm{M}_{3}$. The orthogonal projections
$\mathrm{M}_{3}\longrightarrow\mathrm{M}_{2}\oplus\mathrm{M}_{1}\longrightarrow\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R})$
restrict by Lemma 2.5 to
$\mathcal{D}(\mathrm{M}_{3})\stackrel{{\scriptstyle\pi_{1}}}{{\longrightarrow}}\mathcal{D}(\mathrm{M}_{2}\oplus\mathrm{M}_{1})\stackrel{{\scriptstyle\pi_{2}}}{{\longrightarrow}}\mathcal{D}(\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R}))\,.$
The map $\pi_{1}$ is open by Prop. 6.3. By Lemma 3.7, the map $\pi_{2}$ is the direct convex sum
$\mathrm{Re}\oplus_{\mathrm{c}}\operatorname{id}:\mathcal{D}(\mathrm{M}_{2})\oplus_{\mathrm{c}}\{1\}\to\mathcal{D}(\mathrm{M}_{2}(\mathbb{R}))\oplus_{\mathrm{c}}\{1\}$
of $\mathrm{Re}:\mathcal{D}(\mathrm{M}_{2})\to\mathcal{D}(\mathrm{M}_{2}(\mathbb{R}))$ and the identity map $\operatorname{id}:\{1\}\to\{1\}$. The map $\mathrm{Re}$ is open by Prop. 7.1, hence $\pi_{2}$ is open by Prop. 3.2. The orthogonal projection $\mathcal{D}(\mathrm{M}_{3})\to\mathcal{D}(\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R}))$ is open, as it is a composition of two open maps.
8. Topology simplified by algebra
Thm. 1.2 can simplify topology problems. Given topological spaces $K,L$, a map $f:K\to L$ is continuous [27] at $x\in K$ if the preimage of every neighborhood of $f(x)$ in $L$ is a neighborhood of $x$ in $K$.
Lemma 8.1.
Let $\mathcal{R}_{1},\mathcal{R}_{2}$ be real matrix systems on $\mathbb{C}^{n}$ such that $\mathcal{R}_{2}\subset\mathcal{R}_{1}$. Let $\pi_{1}:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R}_{1})$ and $\pi_{2}:\mathcal{D}(\mathcal{R}_{1})\to\mathcal{D}(\mathcal{R}_{2})$ denote the orthogonal projections and assume the orthogonal projection $\pi_{2}\circ\pi_{1}:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R}_{2})$ is open. Let $f:\mathcal{D}(\mathcal{R}_{2})\to T$ be a map to a topological space $T$. Let $\rho\in\mathcal{D}(\mathcal{R}_{1})$.
  1. a)
    The map $f\circ\pi_{2}:\mathcal{D}(\mathcal{R}_{1})\to T$ is open at $\rho$ if and only if $f:\mathcal{D}(\mathcal{R}_{2})\to T$ is open at $\pi_{2}(\rho)$.
  2. b)
    The map $f\circ\pi_{2}:\mathcal{D}(\mathcal{R}_{1})\to T$ is continuous at $\rho$ if and only if the map $f:\mathcal{D}(\mathcal{R}_{2})\to T$ is continuous at $\pi_{2}(\rho)$.
Proof.
The orthogonal projection $\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R}_{2})$ equals indeed $\pi_{2}\circ\pi_{1}$ by Lemma 2.5.
We begin with the implication $\Rightarrow$” of a). If $\mathcal{N}_{2}\subset\mathcal{D}(\mathcal{R}_{2})$ is a neighborhood of $\pi_{2}(\rho)$, then
$f(\mathcal{N}_{2})=(f\circ\pi_{2})\circ\pi_{2}^{-1}(\mathcal{N}_{2})$
is a neighborhood of $f(\pi_{2}(\rho))$ because $\pi_{2}$ is continuous and $f\circ\pi_{2}$ is open at $\rho$. Regarding the implication $\Leftarrow$”, we choose a neighborhood $\mathcal{N}_{1}\subset\mathcal{D}(\mathcal{R}_{1})$ of $\rho$. Then
$f\circ\pi_{2}(\mathcal{N}_{1})=f\circ(\pi_{2}\circ\pi_{1})\circ\pi_{1}^{-1}(\mathcal{N}_{1})$
is a neighborhood of $f(\pi_{2}(\rho))$ because $\pi_{1}$ is continuous, $\pi_{2}\circ\pi_{1}$ is open, and $f$ is open at $\pi_{2}(\rho)$.
To prove b) we choose a neighborhood $\mathcal{N}_{T}\subset T$ of $f(\pi_{2}(\rho))$. Regarding the implication $\Rightarrow$”, the preimage
$f^{-1}(\mathcal{N}_{T})=(\pi_{2}\circ\pi_{1})\circ\pi_{1}^{-1}\circ(f\circ\pi_{2})^{-1}(\mathcal{N}_{T})$
is a neighborhood of $\pi_{2}(\rho)$, because $f\circ\pi_{2}$ is continuous at $\rho$, the map $\pi_{1}$ is continuous, and $\pi_{2}\circ\pi_{1}$ is open. Regarding the implication $\Leftarrow$”, the preimage
$(f\circ\pi_{2})^{-1}(\mathcal{N}_{T})=\pi_{2}^{-1}\circ f^{-1}(\mathcal{N}_{T})$
is a neighborhood of $\rho$, as $f$ is continuous at $\pi_{2}(\rho)$, and $\pi_{2}$ is continuous. ∎
Remark 8.2 (Simplifying openness problems).
Let $\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R})$ be the orthogonal projection to a real matrix system $\mathcal{R}$ on $\mathbb{C}^{n}$ and let $\mathcal{A}$ be a *-subalgebra of $\mathrm{M}_{n}$ such that $\mathcal{R}\subset\mathcal{A}$. Then $\pi=\pi_{\mathcal{R}}\circ\pi_{\mathcal{A}}$ factors into the orthogonal projections $\pi_{\mathcal{A}}:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{A})$ and $\pi_{\mathcal{R}}:\mathcal{D}(\mathcal{A})\to\mathcal{D}(\mathcal{R})$. For every $\rho\in\mathcal{D}(\mathrm{M}_{n})$ the map $\pi$ is open at $\rho$ if and only if $\pi_{\mathcal{R}}$ is open at $\pi_{\mathcal{A}}(\rho)$.
Indeed, the map $\pi$ factors by Lemma 2.5. The second claim follows from Lemma 8.1 a), by letting $\mathcal{R}_{1}:=\mathrm{M}_{n}$ and $\mathcal{R}_{2}:=\mathcal{A}$, and by taking $\pi_{\mathcal{R}}$ as the map $f:\mathcal{D}(\mathcal{R}_{2})\to T$. The assumptions of the lemma are met since $\pi_{\mathcal{A}}$ is open by Thm. 1.2.
The following examples demonstrate the use of Rem. 8.2. Continuity problems are a topic of Sec. 9 below.
Example 8.3.
Keeping the notation of $\pi=\pi_{\mathcal{R}}\circ\pi_{\mathcal{A}}$ from Rem. 8.2, we take $\mathcal{R}:=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{3},X\oplus 1,Z\oplus 0)$ and $\mathcal{A}:=\mathrm{M}_{2}\oplus\mathrm{M}_{1}$.
  1. a)
    By Ex. 3.5 a), the map $\pi_{\mathcal{R}}$ is not open at any point in the half-open segment $\mathcal{G}_{0}=\big{(}\ket{+}\!\!\bra{+}\oplus 0,0\oplus 1\big{]}\subset\mathcal{D}(\mathcal{A})$ and open at every point in the complement. By Rem. 8.2, the map $\pi:\mathcal{D}(\mathrm{M}_{3})\to\mathcal{D}(\mathcal{R})$ is not open at any point in $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{0})$ and open at every point in the complement.
  2. b)
    To describe $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{0})$, we study the fibers of the orthogonal projection
    $\mathcal{H}(\mathrm{M}_{3})\to\mathcal{H}(\mathcal{A})\,,\quad\begin{pmatrix}A&\ket{\varphi}\\ \bra{\varphi}&c\end{pmatrix}\mapsto\begin{pmatrix}A&0\\ 0&c\end{pmatrix}\,,$
    where $A\in\mathcal{H}(\mathrm{M}_{2})$, |$\ket{\varphi}\in\mathbb{C}^{2}\cong\mathbb{C}^{2\times 1}$, and $c\in\mathbb{R}\cong\mathcal{H}(\mathrm{M}_{1})$. Every point in $\mathcal{G}=\big{[}\ket{+}\!\!\bra{+}\oplus 0,0\oplus 1\big{]}$ is of the form $(1-\lambda)\ket{+}\!\!\bra{+}\oplus\lambda$ for some $\lambda\in[0,1]$. Using the generalized Schur complement (Rem. 6.2), one verifies that the fiber of $\pi_{\mathcal{A}}$ over this point is the set of all matrices
    $\rho(\lambda,z):=\begin{pmatrix}(1-\lambda)\ket{+}\!\!\bra{+}&z\ket{+}\\ \overline{z}\bra{+}&\lambda\end{pmatrix}\,,\quad z\in\mathbb{C}\,,\quad|z|^{2}\leq\lambda(1-\lambda)\,,$
    where $|z|$ denotes the absolute value of $z\in\mathbb{C}$. In conclusion, the orthogonal projection $\pi:\mathcal{D}(\mathrm{M}_{3})\to\mathcal{D}(\mathcal{R})$ is not open at any point of
    $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{0})=\{\rho(\lambda,z)\colon z\in\mathbb{C},|z|^{2}\leq\lambda(1-\lambda),\lambda\in(0,1]\}$
    and open at every point in the complement.
  3. c)
    We verify a claim made in Ex. 4.6. The segment $\mathcal{G}=\big{[}\ket{+}\!\!\bra{+}\oplus 0,0\oplus 1\big{]}$ is the fiber of $\pi_{\mathcal{R}}$ over $M=\frac{1}{2}(\ket{+}\!\!\bra{+}\oplus 1)\in\mathcal{D}(\mathcal{R})$ by Ex. 3.5 b), so $\pi_{\mathcal{A}}^{-1}(\mathcal{G})$ is the fiber of $\pi=\pi_{\mathcal{R}}\circ\pi_{\mathcal{A}}$ over $M$. As recalled in part a) above, the map $\pi$ is open at |++|$\ket{+}\!\!\bra{+}\oplus 0$ but not open at any point of $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{0})$. We now observe that $\pi$ is not open at any point in the relative interior of $\pi^{-1}(M)$, as we have the chain of inclusions
[…]
    $\textstyle=\operatorname{ri}(\pi_{\mathcal{R}}^{-1}(M))=\operatorname{ri}(\mathcal{G})=\mathcal{G}_{0}\setminus\{0\oplus 1\}\subset\mathcal{G}_{0}\,,$
    whose first equality holds by [45, Thm. 6.6].
Example 8.4.
The recipe of Rem. 8.2 helps analyze the openness of the orthogonal projection $\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R}(P,Q))$ to the real matrix system
$\mathcal{R}(P,Q)=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{n},P,Q)$
generated by two orthogonal projections $P,Q\in\mathrm{M}_{n}$, that is to say, matrices satisfying $P=P^{2}=P^{\ast}$ and $Q=Q^{2}=Q^{\ast}$. It is well known that the matrix system $\mathcal{R}(P,Q)$ is included in a surprisingly small *-subalgebra of $\mathrm{M}_{n}$, see Coro. 2.2 in the survey [8] by Böttcher and Spitkovsky, and the references therein. More precisely, there exists a unitary $n\times n$ matrix $U$, nonnegative integers $m_{1}\leq 4$ and $m_{2}$, and positive integers $k_{i}$, $i=1,\ldots,m_{1}$ satisfying $k_{1}+\dots+k_{m_{1}}+2m_{2}=n$, such that $\mathcal{R}:=U\mathcal{R}(P,Q)U^{\ast}$ is included in
$\mathcal{A}:=\big{(}\bigoplus_{i=1}^{m_{1}}\mathrm{M}_{1}\otimes\mathds{1}_{k_{i}}\big{)}\oplus\big{(}\bigoplus_{j=1}^{m_{2}}\mathrm{M}_{2}\big{)}\,.$
Since $\rho\mapsto U\rho U^{\ast}$ is a homeomorphism of $\mathcal{D}(\mathrm{M}_{n})$, the openness problems of the orthogonal projections $\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R}(P,Q))$ and $\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R})$ are equivalent. The second one is substantially simplified by the method of Rem. 8.2 as $\mathcal{D}(\mathcal{A})$ has a rather simple shape. It is the direct convex sum of several three-dimensional Euclidean balls and a simplex of dimension at most three by Lemma 3.4 and Ex. 2.3. This observation should also simplify the strong continuity problem for the numerical range $W(P+\mathrm{i}Q)$.
9. Continuity in quantum information theory
We discuss continuity problems of entropic inference maps and of measures of correlation. We assume that $\mathcal{R}$ is a real matrix system on $\mathbb{C}^{n}$ and that, without loss of generality (see Rem. 2.4), we have $\mathcal{R}\subset\mathcal{H}(\mathrm{M}_{n})$.
Example 9.1 (Maximum entropy inference I).
The purpose of the maximum entropy inference method is to update a prior probability distribution if new information becomes available in the form of constraints that specify a set of possible posterior probability distributions. The preferred posterior is that which minimizes the relative entropy from the prior subject to the available constraints, see Chap. 8 in Caticha’s book [12] and the references therein. An analogous quantum mechanical inference method can be defined by replacing probability distributions with density matrices and the standard relative entropy with the Umegaki relative entropy. The axiomatic foundations of the maximum entropy inference method were settled for probability distributions in the 1980’s, see Chap. 6 in [12, pp. 157–160]. More than 30 years later, the search for axioms of the quantum inference [3] found a convincing answer in the work of Vanslette [51].
Linear constraints [^4] Linear constraints can be defined in terms of expectation values. Let $A_{1},\dots,A_{k}$ be hermitian $n\times n$ matrices such that $\mathcal{R}=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{n},A_{1},\dots,A_{k})$. The observables represented by $A_{1},\dots,A_{k}$ have the expectation values $v(\rho)=(\langle\rho,A_{1}\rangle,\dots,\langle\rho,A_{k}\rangle){}^{\mathrm{T}}$ if $\rho\in\mathcal{D}(\mathrm{M}_{n})$ is the system state [7]. The fiber $\pi^{-1}(\sigma)$ over $\sigma\in\mathcal{D}(\mathcal{R})$ is the set of $\rho\in\mathcal{D}(\mathrm{M}_{n})$ whose expectation values are $v(\rho)=v(\sigma)$, see Lemma 4.1. on $\mathcal{D}(\mathrm{M}_{n})$ are defined by the orthogonal projection
$\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R})\,.$
The relative entropy $S:\mathcal{D}(\mathrm{M}_{n})\times\mathcal{D}(\mathrm{M}_{n})\to[0,+\infty]$ is an asymmetric distance. It is defined by $S(\rho_{1},\rho_{2}):=\operatorname{tr}[\rho_{1}(\log(\rho_{1})-\log(\rho_{2}))]$ if the range of $\rho_{1}$ is included in the range of $\rho_{2}$ and by $S(\rho_{1},\rho_{2}):=+\infty$ otherwise, for all $\rho_{1},\rho_{2}\in\mathcal{D}(\mathrm{M}_{n})$. Let $\tau\in\mathcal{D}(\mathrm{M}_{n})$, the prior, be a density matrix of maximal rank $n$. Then
$\phi_{\tau}:\mathcal{D}(\mathrm{M}_{n})\to\mathbb{R}\,,\quad\rho\mapsto-S(\rho,\tau)$
is continuous and strictly concave. So the maximum entropy inference map
$\Psi_{\tau}:\mathcal{D}(\mathcal{R})\to\mathcal{D}(\mathrm{M}_{n})\,,\quad\sigma\mapsto\operatorname{argmax}_{\rho\in\pi^{-1}(\sigma)}\phi_{\tau}(\rho)$ (9.1)
is well defined, see [53, Def. 1.1] and the references therein. Discontinuities of this inference map [56] aroused interest in theoretical physics [13, 26, 36, 57]. The map (9.1) is continuous for commutative real matrix systems $\mathcal{R}$, for example for the inference of probability distributions mentioned above.
We quote [53, Thm. 4.9].
Theorem 9.2.
Let $\sigma\in\mathcal{D}(\mathcal{R})$. Then $\Psi_{\tau}$ is continuous at $\sigma$ if and only if $\pi$ is open at $\Psi_{\tau}(\sigma)$.
In what follows, continuity and openness problems will be simplified by factoring $\pi=\pi_{\mathcal{R}}\circ\pi_{\mathcal{A}}$ through a *-subalgebra $\mathcal{A}$ of $\mathrm{M}_{n}$ that includes $\mathcal{R}$, using the notation of Rem. 8.2.
Example 9.3 (Maximum entropy inference II).
Before simplifying the openness problem of $\pi$, we describe the set $\Psi_{\tau}(\mathcal{D}(\mathcal{R}))$ of posteriors, as the openness only matters for points in this set by Thm. 9.2. Recalling $\mathcal{R}=\mathcal{H}(\mathcal{R})$, we define
$\mathcal{E}_{\tau}(\mathcal{R}):=\big{\{}\frac{e^{\log(\tau)+A}}{\operatorname{tr}(e^{\log(\tau)+A})}\mid A\in\mathcal{R}\big{\}}\,.$
The manifold $\mathcal{E}_{\tau}(\mathcal{R})$ is known as a Gibbsian family or exponential family, see [56, 38, 53, 43] and the references therein. By (D5) in [53] we have
$\Psi_{\tau}(\mathcal{D}(\mathcal{R}))=\{\rho_{1}\in\mathcal{D}(\mathrm{M}_{n}):\inf_{\rho_{2}\in\mathcal{E}_{\tau}(\mathcal{R})}S(\rho_{1},\rho_{2})=0\}\,.$ (9.2)
The right-hand side of (9.2) is the reverse information closure or rI-closure [18] of $\mathcal{E}_{\tau}(\mathcal{R})$, which is a subset of the Euclidean closure of $\mathcal{E}_{\tau}(\mathcal{R})$.
  1. a)
    If $\tau=\mathds{1}_{n}/n$ is the uniform prior, then
    $\phi_{\tau}(\rho)=-S(\rho,\mathds{1}_{n}/n)=S(\rho)-\log(n)\,,\quad\rho\in\mathcal{D}(\mathrm{M}_{n})$
    is the von Neumann entropy $S(\rho):=-\operatorname{tr}[\rho\log(\rho)]$ up to a constant. By functional calculus, $\mathcal{E}_{\tau}(\mathcal{R})$ is included in $\mathcal{A}$ and so is the set of posteriors $\Psi_{\tau}(\mathcal{D}(\mathcal{R}))$ as per (9.2), because $\mathcal{A}$ is closed. Rem. 8.2 then shows that for every $\rho\in\Psi_{\tau}(\mathcal{D}(\mathcal{R}))$ the openness of $\pi$ at $\rho$ is equivalent to the openness of $\pi_{\mathcal{R}}=\pi|_{\mathcal{D}(\mathcal{A})}$ at $\rho$ (the same conclusion is true for every prior $\tau$ in $\mathcal{A}$). This can simplify the problem if $\mathcal{A}$ has a simpler structure or a smaller dimension than $\mathrm{M}_{n}$. An example is given in Ex. 8.4 above.
  2. b)
    If the prior $\tau$ lies outside of $\mathcal{A}$ then $\mathcal{E}_{\tau}(\mathcal{R})$ is disjoint from $\mathcal{D}(\mathcal{A})$, again by functional calculus. Rem. 8.2 then shows that for every $\rho\in\Psi_{\tau}(\mathcal{D}(\mathcal{R}))$ the openness of $\pi$ at $\rho$ is equivalent to the openness of $\pi_{\mathcal{R}}$ at $\pi_{\mathcal{A}}(\rho)$. This is an even greater simplification than in part a) above, because the analysis of $\pi$ on $\mathcal{D}(\mathrm{M}_{n})$ is reduced to that of $\pi_{\mathcal{R}}$ on $\mathcal{D}(\mathcal{A})$.
  3. c)
    Independently of $\mathcal{A}$, it sometimes helps that the posterior $\Psi_{\tau}(\sigma)$ is contained in the relative interior of the fiber $\pi^{-1}(\sigma)$ over $\sigma$ for every $\sigma\in\mathcal{D}(\mathcal{R})$ and prior $\tau$ by Coro. 5.7 and Lemma 5.8 in [53]. As an example, the orthogonal projection $\pi:\mathcal{D}(\mathrm{M}_{3})\to\mathcal{D}(\mathcal{R})$ to the real matrix system $\mathcal{R}$ in Ex. 8.3 is not open anywhere in the relative interior of the fiber over a certain point $M$ and open at every point in the complement of that fiber. Thm. 9.2 then shows that $\Psi_{\tau}$ is discontinuous at $M$ and continuous everywhere else in the ellipse $\mathcal{D}(\mathcal{R})$ for every prior $\tau$.
From here on, we assume the prior $\tau:=\mathds{1}_{n}/n$ be uniform. Using the von Neumann entropy $S=\phi_{\tau}+\log(n)$, we write the inference map (9.1) from Ex. 9.1 as
$\Psi:\mathcal{D}(\mathcal{R})\to\mathcal{D}(\mathrm{M}_{n})\,,\quad\sigma\mapsto\operatorname{argmax}_{\rho\in\pi^{-1}(\sigma)}S(\rho)\,.$
We also write $\mathcal{E}(\mathcal{R}):=\mathcal{E}_{\tau}(\mathcal{R})$ for the exponential family of Ex. 9.3.
Example 9.4 (Maximum entropy inference III).
An important example from physics is the real matrix system of local Hamiltonians [13, 59].
Every unit $i\in\Omega:=\{1,2,\ldots,N\}$ of an $N$-qubit system is associated with a copy $\mathcal{A}_{i}$ of the algebra $\mathrm{M}_{2}$. The subsystem with units in a subset $\nu\subset\Omega$ is associated with the tensor product algebra $\mathcal{A}_{\nu}:=\bigotimes_{i\in\nu}\mathcal{A}_{i}$, whose identity we denote by $\mathds{1}_{\nu}$. We have $\mathcal{A}_{\Omega}=\mathrm{M}_{n}$ for $n=2^{N}$. The algebra $\mathcal{A}_{\nu}$ embeds into $\mathcal{A}_{\Omega}$ via the map $\mathcal{A}_{\nu}\to\mathcal{A}_{\Omega}$, $A\mapsto A\otimes\mathds{1}_{\bar{\nu}}$, where $\bar{\nu}=\Omega\setminus\nu$ is the complement of $\nu$. Let $\mathfrak{g}$ be a family of subsets of $\Omega$. A $\mathfrak{g}$-local Hamiltonian is a hermitian matrix in $\mathcal{A}_{\Omega}$ of the form
$\sum_{\nu\in\mathfrak{g}}A_{\nu}\otimes\mathds{1}_{\bar{\nu}}\,,\quad A_{\nu}\in\mathcal{H}(\mathcal{A}_{\nu})\,,\quad\nu\in\mathfrak{g}\,.$
We denote the real matrix system of all $\mathfrak{g}$-local Hamiltonians by $\mathcal{R}_{\mathfrak{g}}$ and the orthogonal projection by $\pi_{\mathfrak{g}}:\mathcal{D}(\mathcal{A}_{\Omega})\to\mathcal{D}(\mathcal{R}_{\mathfrak{g}})$.
The partial trace $\operatorname{tr}_{\bar{\nu}}:\mathcal{A}_{\Omega}\to\mathcal{A}_{\nu}$ is the adjoint of the embedding $\mathcal{A}_{\nu}\to\mathcal{A}_{\Omega}$ and satisfies $\langle A\otimes\mathds{1}_{\bar{\nu}},B\rangle=\langle A,\operatorname{tr}_{\bar{\nu}}(B)\rangle$ for every $A\in\mathcal{A}_{\nu}$, $B\in\mathcal{A}_{\Omega}$. The partial trace $\operatorname{tr}_{\bar{\nu}}(\rho)$ of $\rho\in\mathcal{D}(\mathcal{A}_{\Omega})$ is a density matrix of $\mathcal{A}_{\nu}$ called reduced density matrix. Let
$\operatorname{red}_{\mathfrak{g}}:\mathcal{A}_{\Omega}\to\prod_{\nu\in\mathfrak{g}}\mathcal{A}_{\nu}\,,\quad A\mapsto[\operatorname{tr}_{\bar{\nu}}(A)]_{\nu\in\mathfrak{g}}$
denote the map from $\mathcal{A}_{\Omega}$ to the cartesian product of the algebras $(\mathcal{A}_{\nu})_{\nu\in\mathfrak{g}}$ that assigns reduced density matrices.
Linear constraints on $\mathcal{D}(\mathcal{A}_{\Omega})$ have been defined in terms of reduced density matrices, see [38, 13] and [59, Sec. 1.4.2]. This is formalized in the following diagram, which commutes by formula (19) in [55]. (Obvious restrictions of the domain and codomain of $\operatorname{red}_{\mathfrak{g}}$ are omitted in the sequel.)
${\mathcal{D}(\mathcal{A}_{\Omega})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{\mathfrak{g}}}$$\scriptstyle{\operatorname{red}_{\mathfrak{g}}}$${\operatorname{red}_{\mathfrak{g}}[\mathcal{D}(\mathcal{A}_{\Omega})]}$${\mathcal{D}(\mathcal{R}_{\mathfrak{g}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{red}_{\mathfrak{g}}}$
The fiber $\pi_{\mathfrak{g}}^{-1}(\sigma)$ over $\sigma\in\mathcal{D}(\mathcal{R}_{\mathfrak{g}})$ is the set of all $\rho\in\mathcal{D}(\mathcal{A}_{\Omega})$ such that $\operatorname{red}_{\mathfrak{g}}(\rho)=\operatorname{red}_{\mathfrak{g}}(\sigma)$. Thm. 9.2 proves that the pullback $\Xi:=\Psi\circ\operatorname{red}_{\mathfrak{g}}^{-1}$ of the inference map $\Psi$ under $\operatorname{red}_{\mathfrak{g}}^{-1}$ is continuous at $(\rho_{\nu})_{\nu\in\mathfrak{g}}\in\operatorname{red}_{\mathfrak{g}}[\mathcal{D}(\mathcal{A}_{\Omega})]$ if and only if $\operatorname{red}_{\mathfrak{g}}$ is open at $\Xi[(\rho_{\nu})_{\nu\in\mathfrak{g}}]\in\mathcal{D}(\mathcal{A}_{\Omega})$.
Chen et al. [13, Ex. 4] discovered a discontinuity of $\Xi$ for $N=3$ qubits and $\mathfrak{g}:=\{\{1,2\},\{2,3\},\{3,1\}\}$ at $(\rho_{\nu})_{\nu\in\mathfrak{g}}\in\operatorname{red}_{\mathfrak{g}}[\mathcal{D}(\mathcal{A}_{\Omega})]$, where
$\rho_{\nu}:=\frac{1}{2}(\ket{00}\!\!\bra{00}+\ket{11}\!\!\bra{11})\,,\quad\nu\in\mathfrak{g}$
and they offered an interesting interpretation in terms of phase transitions. The map $\operatorname{red}_{\mathfrak{g}}$ being open [^5] The openness of $\operatorname{red}_{\mathfrak{g}}$ at $\rho\in\mathcal{D}(\mathcal{A}_{\Omega})$ is a priori weaker than the continuity of $\Xi$ at $\operatorname{red}_{\mathfrak{g}}(\rho)$. The continuity means that any sufficiently small change of $\operatorname{red}_{\mathfrak{g}}(\rho)$ can be matched by an arbitrarily small change of $\rho$ inside the image of $\Xi$ and not just anywhere in $\mathcal{D}(\mathcal{A}_{\Omega})$. Somewhat surprisingly, the two propositions are equivalent by Thm. 9.2. at $\rho\in\mathcal{D}(\mathcal{A}_{\Omega})$ means that any sufficiently small change of $\operatorname{red}_{\mathfrak{g}}(\rho)$ in $\operatorname{red}_{\mathfrak{g}}[\mathcal{D}(\mathcal{A}_{\Omega})]$ is matched by an arbitrarily small change of $\rho$ within $\mathcal{D}(\mathcal{A}_{\Omega})$. Conversely, if the openness fails, then there are arbitrarily small changes of $\operatorname{red}_{\mathfrak{g}}(\rho)$ that can only be matched by changes of $\rho$ beyond some strictly positive threshold (in the metric sense). Loosely speaking, a small change of a subsystem abruptly changes the entire system. Such behavior is associated with phase transitions. It motivates every attempt to study the openness of $\pi_{\mathfrak{g}}$. This should be done by tuning the interaction pattern $\mathfrak{g}$ to a concrete system. Whether enclosing $\mathcal{R}_{\mathfrak{g}}$ into a *-subalgebra of $\mathcal{A}_{\Omega}$ could simplify this problem, as suggested by Exa. 9.3, is not yet clarified.
We finish this paper with a map whose continuity is more subtle than that of the inference.
Example 9.5 (Entropy distance I).
The entropy distance from the exponential family $\mathcal{E}(\mathcal{R})$ is defined by
$d:\mathcal{D}(\mathrm{M}_{n})\to\mathbb{R}\,,\quad\rho_{1}\mapsto\inf_{\rho_{2}\in\mathcal{E}(\mathcal{R})}S(\rho_{1},\rho_{2})$ (9.3)
and equals the difference
$d(\rho)=S(\Psi\circ\pi(\rho))-S(\rho)\,,\quad\rho\in\mathcal{D}(\mathrm{M}_{n})$ (9.4)
between the value of the von Neumann entropy at $\rho$ and the maximal value on the fiber of $\pi$ that contains $\rho$, see p. 1288 in [53]. Formula (9.4) suggests studying the continuity of $d$ through the rI-projection
$\Pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathrm{M}_{n})\,,\quad\Pi:=\Psi\circ\pi\,.$
As per Def. 5.2 and equation (D8) in [53], the density matrix $\Pi(\rho_{1})$ is the generalized rI-projection of $\rho_{1}\in\mathcal{D}(\mathrm{M}_{n})$ to $\mathcal{E}(\mathcal{R})$, which is a well-known concept in probability theory [18], and which is defined as follows. A sequence $(\tau_{i})\subset\mathcal{D}(\mathrm{M}_{n})$ rI-converges to $\rho_{2}\in\mathcal{D}(\mathrm{M}_{n})$ if $\lim_{i}S(\rho_{2},\tau_{i})=0$ holds. If every sequence $(\tau_{i})\subset\mathcal{E}(\mathcal{R})$ satisfying $\lim_{i}S(\rho_{1},\tau_{i})=d(\rho_{1})$ rI-converges, independently of the sequence, to a unique $\rho_{2}\in\mathcal{D}(\mathrm{M}_{n})$, not necessarily in $\mathcal{E}(\mathcal{R})$, then $\rho_{2}$ is the generalized rI-projection of $\rho_{1}$ to $\mathcal{E}(\mathcal{R})$.
We quote from Lemma 5.15 and Lemma 4.5 in [53].
Lemma 9.6.
  1. a)
    For every $\rho\in\mathcal{D}(\mathrm{M}_{n})$ the rI-projection $\Pi$ is continuous at $\rho$ if and only if the entropy distance $d$ is continuous at $\rho$.
  2. b)
    For every $\sigma\in\mathcal{D}(\mathcal{R})$, the inference map $\Psi$ is continuous at $\sigma$ if and only if $d$ is continuous at every point in the fiber $\pi^{-1}(\sigma)$.
Example 9.7 (Entropy distance II).
The entropy distance from the exponential family $\mathcal{E}(\mathcal{R}_{\mathfrak{g}})$ of local Hamiltonians (Ex. 9.4) is interesting because it quantifies many-body correlations. Amari [4] and Ay [6] studied this type of correlation measures in probability theory. Linden et al. [33] introduced it to quantum mechanics as a difference of von Neumann entropies like formula (9.4), see also [59, Sec. 1.4.2]. Zhou [60] proved the equality of the two representations (9.3) and (9.4) for density matrices of maximal rank $n$, see also [38]; the equality is true without rank restrictions as stated in Ex. 9.5.
Proposition 9.8.
For every $\rho\in\mathcal{D}(\mathrm{M}_{n})$, the entropy distance $d$ is continuous at $\rho$ if and only if its restriction $d|_{\mathcal{D}(\mathcal{A})}$ is continuous at $\pi_{\mathcal{A}}(\rho)$.
Proof.
Let $\rho\in\mathcal{D}(\mathrm{M}_{n})$. By Lemma 9.6 a), the entropy distance $d$ is continuous at $\rho$ if and only if the rI-projection $\Pi$ is continuous at $\rho$. The inference map
$\Psi^{\mathcal{A}}:\mathcal{D}(\mathcal{R})\to\mathcal{D}(\mathcal{A})\,,\quad\sigma\mapsto\operatorname{argmax}_{\eta\in\pi_{\mathcal{R}}^{-1}(\sigma)}S(\eta)$
has the same values as $\Psi$, whose image $\Psi(\mathcal{D}(\mathcal{R}))$ is included in $\mathcal{D}(\mathcal{A})$ by Ex. 9.3 a). Therefore,
$\Pi=\Psi\circ\pi=\Psi\circ\pi_{\mathcal{R}}\circ\pi_{\mathcal{A}}$
is continuous at $\rho$ if and only if $\Psi^{\mathcal{A}}\circ\pi_{\mathcal{R}}\circ\pi_{\mathcal{A}}$ is continuous at $\rho$, if and only if $\Pi^{\mathcal{A}}\circ\pi_{\mathcal{A}}$ is continuous at $\rho$, where
$\Pi^{\mathcal{A}}:\mathcal{D}(\mathcal{A})\to\mathcal{D}(\mathcal{A})\,,\quad\Pi^{\mathcal{A}}:=\Psi^{\mathcal{A}}\circ\pi_{\mathcal{R}}\,.$
Since $\pi_{\mathcal{A}}$ is open by Thm. 1.2, it follows from Lemma 8.1 b) that $\Pi^{\mathcal{A}}\circ\pi_{\mathcal{A}}$ is continuous at $\rho$ if and only if $\Pi^{\mathcal{A}}$ is continuous at $\pi_{\mathcal{A}}(\rho)$. By equation (9.4),
$d|_{\mathcal{D}(\mathcal{A})}(\eta)=S(\Psi^{\mathcal{A}}\circ\pi_{\mathcal{R}}(\eta))-S(\eta)\,,\quad\eta\in\mathcal{D}(\mathcal{A})$
holds. Hence, $\Pi^{\mathcal{A}}$ is continuous at $\pi_{\mathcal{A}}(\rho)$ if and only if $d|_{\mathcal{D}(\mathcal{A})}$ is continuous at $\pi_{\mathcal{A}}(\rho)$, again by Lemma 5.15 1) in [53]. ∎
Example 9.9 (Entropy distance III).
Prop. 9.8 helps solve the continuity problem of the entropy distance from $\mathcal{E}(\mathcal{R})$ for $\mathcal{R}:=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{3},X\oplus 1,Z\oplus 0)$ using the solution of the continuity problem of the restriction $d|_{\mathcal{D}(\mathcal{A})}$ to the real *-subalgebra $\mathcal{A}:=\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R})$ of $\mathrm{M}_{3}$. This solution was obtained from asymptotic curvature estimates [53]. Real *-subalgebras are excluded from Proposition 9.8 but the conclusion is still true, as Ex. 7.2 can replace Thm. 1.2 in the proof of Prop. 9.8.
We recall from Ex. 3.6 that the generatrix $\mathcal{G}=[\ket{+}\!\!\bra{+}\oplus 0,0\oplus 1]$ of the cone $\mathcal{D}(\mathcal{A})$ is the fiber of the orthogonal projection $\pi_{\mathcal{R}}:\mathcal{D}(\mathcal{A})\to\mathcal{D}(\mathcal{R})$ over $M:=\frac{1}{2}(\ket{+}\!\!\bra{+}\oplus 1)$. Thm. 5.18 in [53] shows that $d|_{\mathcal{D}(\mathcal{A})}$ is discontinuous at every point in the half-open segment
$\mathcal{G}_{d}:=\big{[}\ket{+}\!\!\bra{+}\oplus 0,M\big{)}=\big{\{}(1-\lambda)\ket{+}\!\!\bra{+}\oplus\lambda\colon\lambda\in[0,\frac{1}{2})\big{\}}$
and continuous at every point in the complement [^6] Unitary similarity with respect to $U:=\exp(\mathrm{i}\frac{\pi}{3\sqrt{3}}(X+Y+Z))$ permutes the Pauli matrices cyclicly. With respect to $U\oplus 1$ it matches $\mathcal{R}$ and the problem considered in [53].. Prop. 9.8 then shows that the entropy distance $d$ is discontinuous at every point in $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{d})$ and continuous at every point in the complement.
This result is consistent with the assertion of Ex. 9.3 c) that the inference map $\Psi$ is discontinuous at $M$ and continuous anywhere else in the ellipse $\mathcal{D}(\mathcal{R})$, as required by Lemma 9.6 b). The points in the set $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{d})$ are explicitly described in Ex. 8.3 b).
Acknowledgments. The author is grateful to Ariel Caticha, Achim Clausing, Chi-Kwong Li, Tim Netzer, Ilya M. Spitkovsky, and Ivan Todorov for remarks, feedback, and discussions. The author gave a lecture on the results of this paper in the special session “Numerical Ranges” of the conference IWOTA 2024 in Canterbury, UK, and wishes to thank the organizers for this opportunity.